### v5

# Skew-adjoint linear relations in Banach space

Abstract.
In this paper, we prove the stability theorems for the isotropic perturbations of maximal isotropic subspaces in symplectic Banach spaces. Then we prove a stability theorem for the mod $2$ dimensions of kernel of skew-adjoint linear Fredholm relations between real Banach spaces with index 00. Finally we gives the exactly two path components of the set of skew-adjoint linear Fredholm relations between real Banach spaces with indices 00.
Key words and phrases:
Gap between closed linear subspaces, symplectic Banach spaces, skew-adjoint linear relations
2010 Mathematics Subject Classification:
Primary 53D12; Secondary 58J30
Corresponding author: C. Zhu [zhucf@nankai.edu.cn]
Chaofeng Zhu is supported by National Key R&D Program of China (2020YFA0713300), NSFC, China Grants (11971245, 11771331), Nankai Zhide Foundation and Nankai University.
1. Introduction
M. F. Atiyah and I. M. Singer in [2] considered the set of all skew-adjoint Fredholm operators ${\mathcal{F}}^{1}(X)$ on a real separable Hilbert space $X$ and showed that it is a classifying space for (Atiyah’s) real K-theory. As a consequence, the homotopy groups of ${\mathcal{F}}^{1}(X)$ are $8$-periodic, with a fundamental group that is cyclic of order $2$. A.L. Carey, J. Phillips, H. Schulz-Baldes in [6] constructed a homotopy invariant for paths in ${\mathcal{F}}^{1}(X)$ with invertible endpoints which provides an explicit isomorphism between $\pi_{1}({\mathcal{F}}^{1}(X))$ and ${\mathbb{Z}}_{2}$. The authors call it ${\mathbb{Z}}_{2}$-valued spectral flow, as formally it shares properties with the (${\mathbb{Z}}$-valued) spectral flow for paths of selfadjoint Fredholm operators [see M. F. Atiyah, V. K. Patodi and I. M. Singer[1]]. Recently, N. Doll[9] defined the oriented flow for paths in ${\mathcal{F}}^{1}(X)$ with odd dimensional kernels.
In the Hilbert space setting, one of the most powerful tool of studing selfadjoint operators and skew-adjoint operators is the spectral theory. By the idea of symplectic geometry, B. Booss and the second author [5, Definition 4.2.1] was able to generalize the notion of selfadjoint operators and skew-adjoint operators to the linear relations between two real (complex) linear vector spaces with a nondegenerate sesquilinear form. In the Banach space setting, the classical way of studying selfadjoint operators and skew-adjoint operators fails. Our paper is an attempt to the study of selfadjoint linear relations and skew-adjoint linear relations between Banach spaces. Then it is possible to define the oriented flow in the Banach space situation.
Since the selfadjoint operators form the most important class of operators that appear in applications, the stability of selfadjointness is one of our important problems. In [13, Section V.4], T. Kato had given a nice introduction to this topic. Let $X$ be a Hilbert space. A linear operator $A\colon X\supset\operatorname{dom}(A)\to X$ is symmetric (or selfadjoint) if and only if the graph $\operatorname{graph}(A)$ of $A$ is a isotropic (or Lagrangian) linear subspace of $X\times X$.
In this paper, we shall generalize the theory to the symplectic setting. Note that the proofs in [13, Section V.4] do not apply to the symplectic case.
Throughout this paper, we denote by ${\mathbb{K}}$ the field of real numbers or complex numbers. We denote by ${\mathbb{N}}$, ${\mathbb{Z}}$, ${\mathbb{R}}$ and ${\mathbb{C}}$ the sets of all natural, integral, real and complex numbers respectively. By $S^{1}$ we denote the unit circle in the complex plane. We denote by $\mathrm{i}$ the imaginary unit. Denote by $I_{X}$ the identity map on a set $X$. If there is no confusion, we will omit the subindex $X$. For a Banach space $X$ over ${\mathbb{K}}$, we denote by ${\mathcal{B}}(X)$ the set of bounded linear operators. As in [13, Sectioin III.1.4], the adjoint space $X^{*}$ of a Banach space is defined to be the set of bounded semilinear forms on $X$. We equip the the set of closed linear subspaces ${\mathcal{S}}(X)$ with gap distance $\hat{\delta}$ ([13, Section IV.2.1]). Let $X,Y$ be two Banach spaces. A bounded sesquilinear form $\Omega\colon X\times Y\to{\mathbb{K}}$ define two bounded linear maps $L_{\Omega}\colon X\to Y^{*}$ and $R_{\Omega}\colon Y\to X^{*}$ by
(1) $\Omega(x,y)=(L_{\Omega}x)y=\overline{(R_{\Omega}y)x}\quad\text{ for all }x\in X\text{ and }y\in Y.$
Let $(X,\omega)$ be a symplectic vector space over ${\mathbb{K}}$ with a maximal isotropic subspace. Then $h_{\lambda}$ denotes the sign of $\lambda$ defined by Definition 2.18 below. By definition, a maximal isotropic subspace $\lambda$ is Lagrangian if and only if $h_{\lambda}=0$.
We have the following stability results of maximal isotropic subspaces in symplectic Banach spaces.
Theorem 1.1.
Let $(X,\omega_{0})$ be a symplectic Banach space with a maximal isotropic subspace $\lambda$. We assume that $\gamma_{\lambda}>0$ holds if $h_{\lambda}\neq 0$. Let $(X,\omega)$ be a symplectic Banach space with a closed isotropic subspace of $\mu$. We set
$l(\mu,\omega)\ :$ $=\left(|h_{\lambda}|\gamma_{\lambda}^{-1}\left(\|L_{\omega}-L_{\omega_{0}}\|+\|L_{\omega_{0}}\|\delta(\mu^{\omega},\lambda^{\omega_{0}})(2+\delta(\mu^{\omega},\lambda^{\omega_{0}}))\right)\right)^{\frac{1}{2}}.$
Here $h_{\lambda}$ and $\gamma_{\lambda}$ are defined by Definition 2.18 below, and $l(\mu,\omega)=0$ if $h_{\lambda}=0$.
Assume that there hold that
(2) $\delta(\mu^{\omega},\lambda^{\omega_{0}})+l(\mu,\omega)<\frac{1-\delta(\lambda,\mu)}{1+\delta(\lambda,\mu)}.$
Then $\mu$ is a maximal isotropic subspace of $(X,\omega)$ with $h_{\mu}\in\{h_{\lambda},0\}$.
We can use [11, Lemma 1] to prove Theorem 1.1 if $h_{\lambda}=0$. To prove the general case, we need to study the augmented Morse index for the perturbation of a positive semidefinite symmetric form with possible different domain in a fixed Banach space. Note that in general one can not make the domains fixed.
By the study of the Morse index for the perturbation of a positive semidefinite symmetric form and Theorem 1.1, we have the following result.
Theorem 1.2.
Let $X$ be a Banach space with a continuous family of symplectic structure $\{\omega(s)\}_{s\in J}$, where $J$ is a connected topological space. Let $\{\lambda(s)\}_{s\in J}$ be a continuous family of closed subspace of $X$ such that $\lambda(s)\subset\lambda(s)^{\omega(s)}$ holds for each $s\in J$. Assume that $\{\lambda(s)^{\omega(s)}\}_{s\in J}$ and $\{\lambda(s)^{\omega(s)\omega(s)}\}_{s\in J}$ are also continuous families of closed subspace of $X$. Assume that there is an $s_{0}\in J$ such that $\lambda(s_{0})$ is a maximal isotropic subspace of $(X,\omega(s_{0}))$ satisfying $\gamma_{\lambda(s_{0})}>0$ if $h_{\lambda(s_{0})}\neq 0$, and each $s\in J$ such that $\lambda(s)$ is a maximal isotropic subspace of $(X,\omega(s))$ satisfying $\gamma_{\lambda(s)}>0$ if $h_{\lambda(s)}\neq 0$. Then $\lambda(s)$ is a maximal isotropic subspace of $(X,\omega(s))$ and $h_{\lambda(s)}=h_{\lambda(s_{0})}$ holds for each $s\in J$.
In strong symplectic Banach spaces, some conditions in stability theorems of maximal isotropic subspaces hold. Then these stability theorems can be simplified(see Section 2.5 below).
In 2023, J.M. F. Castillo, W. Cuellar M. González and R. Pino [7, Proposition 8.11] showed that a strong real symplectic Hilbert space $(X,\omega)$ can not be a codimension $1$ linear subspace of a strong real symplectic Hilbert space $(Y,\tilde{\omega})$ with compact $L_{\tilde{\omega}|_{X}}-L_{\omega}$. After their proof, they wrote:
”We do not know whether Proposition 8.11 holds for general Banach spaces.”
Note that [7, Proposition 8.11] is a special case of the stability theorem for the mod 2 dimensions of kernel of bounded skew-adjoint operators of M. F. Atiyah and I. M. Singer [2, p.6] in real Hilbert spaces. In this paper, we shall solve the problem and generalize it to a stability theorem for the mod 2 dimensions of kernel and indeterminacy in real Banach spaces. Here we require that $L_{\Omega}$ is an injective bounded operator only.
Theorem 1.3.
Let $X$ and $Y$ be two real Banach space, and $\Omega_{0}\colon X\times Y\to{\mathbb{R}}$ be a bounded nondegenerate bilinear form.
Let $T:X\leadsto Y$ be a skew-adjoint linear Fredholm relation associated with $\Omega_{0}$ with index 00. Then there is a $\delta>0$ such that for each skew symmetric closed linear relation $S:X\leadsto Y$ with respect to a bounded nondegenerate bilinear form $\Omega$ with $\hat{\delta}(S,T)+\|L_{\Omega}-L_{\Omega_{0}}\|<\delta$, there hold that $S$ is skew-adjoint linear Fredholm relation with index 00, and
(3) $\dim\ker S=\dim\ker T\mod 2.$
The path components of the set of skew-adjoint Fredholm relations between real Banach spaces with index 00 is given as follows.
Theorem 1.4.
Let $X$ and $Y$ be two real Banach space, and $\Omega\colon X\times Y\to{\mathbb{R}}$ be a bounded nondegenerate bilinear form.
Let ${\mathcal{F}}{\mathcal{R}}_{0}^{sk}(\Omega)$ be the set of real skew-adjoint linear Fredholm relations with index 00 with respect to $\Omega$. For $a\in\{0,1\}$, we define
(4) ${\mathcal{F}}{\mathcal{R}}_{a;0}^{sk}(\Omega)\ :$ =$=\{T\in{\mathcal{F}}{\mathcal{R}}_{0}^{sk}(\Omega);\;\dim\ker T=a\mod 2\}.$
Then the space ${\mathcal{R}}_{0}^{sk}(\Omega)$ has exactly two path components ${\mathcal{F}}{\mathcal{R}}_{0;0}^{sk}(\Omega)$ and ${\mathcal{F}}{\mathcal{R}}_{1;0}^{sk}(\Omega)$.
In the finite dimensional case, we can always give a Hilbert structure on $X$. By Proposition 4.1 below, there is a homoemorphism $\varphi_{\Omega}\colon\mathrm{O}(X)\to{\mathcal{F}}{\mathcal{R}}_{0}^{sk}(\Omega)={\mathcal{R}}^{sk}(\Omega)$. Moreover, we have
$\ker(T)=\ker(\varphi_{\Omega}^{-1}(T)-I_{X})$
for each $T\in{\mathcal{R}}^{sk}(\Omega)$. Thus Theorem 1.3 and Theorem 1.4 hold in this case.
In the general Banach space situation, we use the symplectic reduction method and reduce the problems to the finite dimensional case.
The paper is organized as follows. In Section 1, we explain the motivation of the research and state the main results of the paper. In Section 2, we prove Theorem 1.1 and Theorem 1.2. In Section 3, we study the stability theorems with relative bounded perturbations. In Section 4, we prove Theorem 1.3 and Theorem 1.4.
We would like to thank the referees of this paper for their critical reading and very helpful comments and suggestions.
2. Stability theorems in symplectic Banach space
2.1. Fredhom pairs of closed linear subspaces
Let $X$ be a Banach space. We denote by ${\mathcal{S}}(X)$ the set of closed linear subspaces of $X$ and ${\mathcal{S}}^{c}(X)$ the set of complemented closed linear subspaces of $X$ respectively. Let $M,N$ be two closed linear subspaces of $X$, i.e., $M,N\in{\mathcal{S}}(X)$. Denote by $\operatorname{S}_{M}$ the unit sphere of $M$. We recall three common definitions of distances in ${\mathcal{S}}(X)$ (see also [13, Sections IV.2.1 and IV.4.1]):
  • the Hausdorff metric $\hat{d}$;
  • the aperture (gap distance) $\hat{\delta}$, that is not a metric since it does not in general satisfy the triangle inequality, but defines the same topology as the metric $\hat{d}$, called gap topology, and is easier to estimate than $\hat{d}$; and
  • the angular distance (minimum gap) $\widehat{\gamma}$, that is useful in our estimates, though not defining any suitable topology.
Definition 2.1 (The gap between subspaces).
(a) We set
$\hat{d}(M,N)\ =\ d(\operatorname{S}_{M},\operatorname{S}_{N})\\ :=\begin{cases}\max\left\{\begin{matrix}\sup\limits_{u\in\operatorname{S}_{M}}\operatorname{dist}(u,\operatorname{S}_{N}),\\ \sup\limits_{u\in\operatorname{S}_{N}}\operatorname{dist}(u,\operatorname{S}_{M})\end{matrix}\right\},&\text{ if both $M\neq 0$ and $N\neq 0$},\\ 0,&\text{ if $M=N=0$},\\ 2,&\text{ if either $M=0$ and $N\neq 0$ or vice versa}.\end{cases}$
(b) We set
$\delta(M,N)$ ::=:= $\ \begin{cases}\sup\limits_{u\in\operatorname{S}_{M}}\operatorname{dist}(u,N),&\text{if $M\neq\{0\}$},\\ 0,&\text{if $M=\{0\}$},\end{cases}$
$\hat{\delta}(M,N)$ ::=:= $\ \max\{\delta(M,N),\delta(N,M)\}.$
$\hat{\delta}(M,N)$ is called the gap between $M$ and $N$.
(c) We set
$\gamma(M,N)\ :=$$\ \begin{cases}\inf\limits_{u\in M\setminus N}\frac{\operatorname{dist}(u,N)}{\operatorname{dist}(u,M\cap N)}\ (\leq 1),&\text{if $M\not\subset N$},\\ 1,&\text{if $M\subset N$},\end{cases}$
$\hat{\gamma}(M,N)\ :=$ $\ \min\{\gamma(M,N),\gamma(N,M)\}.$
$\hat{\gamma}(M,N)$ is called the minimum gap between $M$ and $N$. If $M\cap N=\{0\}$, we have
$\gamma(M,N)\ =\ \inf\limits_{u\in\operatorname{S}_{M}}\operatorname{dist}(u,N).$
In this paper we shall impose the gap topology on the space ${\mathcal{S}}(X)$ of all closed linear subspaces of a Banach space $X$ and its subset ${\mathcal{S}}^{c}(X)$ of complemented subspaces.
Definition 2.2.
(a) The space of (algebraic) Fredholm pairs of linear subspaces of a vector space $X$ is defined by
(5) ${\mathcal{F}}^{2}_{\operatorname{alg}}(X)\ :=\{(M,N);\;\dim(M\cap N)<+\infty,\;\dim X/(M+N)<+\infty\}$
with
(6) $\operatorname{index}(M,N;X)=\operatorname{index}(M,N)\ :=\dim(M\cap N)-\dim X/(M+N).$
(b) In a Banach space $X$, the space of (topological) Fredholm pairs is defined by
(7) ${\mathcal{F}}^{2}(X)\ :=\{(M,N)\in{\mathcal{F}}^{2}_{\operatorname{alg}}(X);\;M,N,\mbox{\ \rm and }M+N\subset X\text{ closed}\}.$
A pair $(M,N)$ of closed subspaces is called semi-Fredholm if $M+N$ is closed, and at least one of $\dim(M\cap N)$ and $\dim\left(X/(M+N)\right)$ is finite.
(c) Let $X$ be a Banach space, $M\in{\mathcal{S}}(X)$ and $k\in{\mathbb{Z}}$. We define
(8) ${\mathcal{F}}_{M}(X)\ :$ =$=\{N\in{\mathcal{S}}(X);\;(M,N)\in{\mathcal{F}}^{2}(X)\},$
(9) ${\mathcal{F}}_{k,M}(X)\ :$ =$=\{N\in{\mathcal{S}}(X);\;(M,N)\in{\mathcal{F}}^{2}(X),\operatorname{index}(M,N)=k\}.$
2.2. Symplectic Banach spaces
Firstly we recall the basic concepts and properties of symplectic functional analysis.
Let $X$ be a vector space over ${\mathbb{K}}$, and $\Omega\colon X\times X\to{\mathbb{K}}$ be a symmtric sesquilinear form or a skew symmtric sesquilinear form. The annihilator of a subspace $\lambda$ of $X$ is defined by
(10) $\lambda^{\Omega}\ :=\{y\in X;\;\Omega(x,y)=0,\;\forall x\in\lambda\}.$
The form $\Omega$ is called non-degenerate, if $X^{\Omega}=\{0\}$.
Definition 2.3.
Let $X$ be a vector space over ${\mathbb{K}}$.
(a) A mapping
$\omega\colon X\times X\longrightarrow{\mathbb{K}}$
is called a symplectic form on $X$, if it is a non-degenerate skew symmtric sesquilinear form.
Then we call $(X,\omega)$ a symplectic vector space.
(b) Let $X$ be a Banach space and $(X,\omega)$ a symplectic vector space. $(X,\omega)$ is called (weak) symplectic Banach space, if $\omega$ is bounded, i.e., $|\omega(x,y)|\leq C\|x\|\|y\|$ for all $x,y\in X$. $(X,\omega)$ is called strong symplectic Banach space, if $\omega$ is bounded and $L_{\omega}^{-1}\in{\mathcal{B}}(X^{*},X)$.
(c) Let $(X_{j},\omega_{j})$ be two symplectic vector spaces. A linear map $L\colon X_{1}\to X_{2}$ is called symplectic if $L^{*}\omega_{2}=\omega_{1}$, i.e., $\omega_{2}(Lx,Ly)=\omega_{1}(x,y)$ for each $x,y\in X_{1}$. The symplectic group $\operatorname{Sp}(X,\omega)$ consists of all surjective symplectic linear maps of a symplectic vector space $(X,\omega)$. We write $\operatorname{Sp}(X)\ :=\operatorname{Sp}(X,\omega)$ if there is no confusion.
(c) A subspace ${\lambda}$ is called symplectic, isotropic, co-isotropic, or Lagrangian if
$\lambda\cap{\lambda}^{\omega}\ =\ \{0\}\,,\quad{\lambda}\,\subset\,{\lambda}^{\omega}\,,\quad{\lambda}\,\supset\,{\lambda}^{\omega}\,,\quad{\lambda}\,\ =\ \,{\lambda}^{\omega}\,,$
respectively. An isotropic subspace ${\lambda}$ is called maximal isotropic, if each isotropic subspace $\mu\supset\lambda$ satisfies $\mu=\lambda$.
(d) The Lagrangian Grassmannian ${\mathcal{L}}(X,\omega)$ consists of all Lagrangian subspaces of $(X,\omega)$. We write ${\mathcal{L}}(X)\ :={\mathcal{L}}(X,\omega)$ if there is no confusion.
(d) The complemented Lagrangian Grassmannian ${\mathcal{C}}{\mathcal{L}}(X,\omega)$ is defined by
(11) ${\mathcal{C}}{\mathcal{L}}(X,\omega)\ :=\{\lambda\in{\mathcal{L}}(X,\omega);\;X=\lambda\oplus\mu\text{ for some }\mu\in{\mathcal{L}}(X,\omega)\}.$
We write ${\mathcal{C}}{\mathcal{L}}(X)\ :={\mathcal{C}}{\mathcal{L}}(X,\omega)$ if there is no confusion.
For a symplectic Banach space $(X,\omega)$, the induced operator $L_{\omega}\colon X\to X^{*}$ defined by (1) is skew-adjoint in the sense that $L_{\omega}^{*}|_{X}=-L_{\omega}$.
A linear relation $A\colon X\rightsquigarrow Y$ between two linear space $X$ and $Y$ is a linear subspace of $X\times Y$. For two linear relatioins $A\colon X\rightsquigarrow Y$ and $B\colon Y\rightsquigarrow Z$, the composite $B\circ A$ is defined by
(12) $B\circ A=\{(x,z)\in X\times Z;\ (x,y)\in A\text{ and }(y,z)\in B\text{ for some }y\in Y\}.$
A linear operator can be viewed as a linear relation by taking its graph. For details, we refer to R. Cross [8] and G. W. Whitehead [15, B.2].
Let $X$ be a Banach space and $M\in{\mathcal{S}}(X)$ be a closed subspace. We define
(13) $M^{\bot}\ :=\{f\in X^{*};\;f(x)=0,\text{ for all }x\in M\}.$
Let $X$ and $Y$ be two vector spaces over ${\mathbb{K}}$. Given a nondegenerate sesquilinear form $\Omega\colon X\times Y\to{\mathbb{K}}$, there is a natural symplectic structure $\omega$ on $Z\ :=X\times Y$ defined by (cf. [3], [5, Definition 4.2.1], [10])
(14) $\omega((x_{1},y_{1}),(x_{2},y_{2}))\ :=\Omega(x_{1},y_{2})-\overline{\Omega(x_{2},y_{1})}$
for all $x_{1},x_{2}\in X$ and $y_{1},y_{2}\in Y$. For each subsets $\lambda$ of $X$ and $\mu$ of $Y$, the annihilator $\lambda^{\Omega,r}$ and $\mu^{\Omega,l}$ (cf. [5, Definition 1.1.1.a]) are defined by
(15) $\lambda^{\Omega,r}\ :=$$\{y\in Y;\ \Omega(x,y)=0\text{ for all }x\in\lambda\},$
(16) $\mu^{\Omega,l}\ :=$$\{x\in X;\ \Omega(x,y)=0\text{ for all }y\in\mu\}.$
Then there are two nature Lagrangian subspaces $X\times\{0\}$ and $\{0\}\times Y$ of $(Z,\omega)$. If $X,Y$ are Banach spaces and $\Omega$ is bounded, by (1), the induced operator of $\omega$ is given by
(19) $L_{\omega}=\left(\begin{array}[]{cc}0&-R_{\Omega}\\ L_{\Omega}&0\end{array}\right).$
Let $X$ be locally convex space. We define $\Omega_{X}\colon X\times X^{*}\to{\mathbb{K}}$ by $\Omega_{X}(x,f)=\overline{f(x)}$ for all $x\in X$ and $f\in X^{*}$. By Hahn-Banach theorem, $\Omega_{X}$ is nondegenerate and continuous. Then $(X\times X^{*},\omega)$ is a symplectic locally convex space.
Definition 2.4.
(cf. [5, Definition 4.2.1]) Let $X$ and $Y$ be two vector spaces over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{K}}$ be a nondegenerate sesquilinear form and $h\in{\mathbb{K}}$ be a number. Let $\omega$ be the associated symplectic form of $\Omega$ defined by (14).
  • (a)
    Let $A\colon X\leadsto Y$ be a linear relation. The $\Omega$-adjoint of $A$ is defined to be $A^{\Omega}\ :=A^{\omega}$. We call $A$ to be $h$-symmetric with respect to $\Omega$ if $hA\subset A^{\Omega}$, and $h$-selfadjoint with respect to $\Omega$ if $hA=A^{\Omega}$, respectively. If $h=1$, we call $A$ symmetric with respect to $\Omega$ and selfadjoint with respect to $\Omega$ respectively. If $h=-1$, we call $A$ skew symmetric with respect to $\Omega$ and skew-adjoint with respect to $\Omega$ respectively. We call $A$ a maximal (skew) symmetric linear relation with respect to $\Omega$, if $A$ is (skew) symmetric linear relation with respect to $\Omega$ and maximal in the set of all (skew) symmetric linear relations with respect to $\Omega$.
  • (b)
    We denote by ${\mathcal{R}}^{sa}(\Omega)$ the set of selfadjoint relations between $X$ and $Y$ with respect to $\Omega$, and ${\mathcal{R}}^{sk}(\Omega)$ the set of skew-adjoint relations between $X$ and $Y$ with respect to $\Omega$, respectively. We denote by ${\mathcal{R}}^{sa}(X)\ :={\mathcal{R}}^{sa}(\Omega_{X})$ and ${\mathcal{R}}^{sk}(X)\ :={\mathcal{R}}^{sk}(\Omega_{X})$ respectively. For $k\in{\mathbb{Z}}$, we define
    ${\mathcal{F}}{\mathcal{R}}^{sa}(\Omega)\ :$ =$=\{A\in{\mathcal{R}}^{sa}(\Omega);\;A\text{ is Fredholm}\},$
    ${\mathcal{F}}{\mathcal{R}}^{sk}(\Omega)\ :$ =$=\{A\in{\mathcal{R}}^{sk}(\Omega);\;A\text{ is Fredholm}\},$
    ${\mathcal{F}}{\mathcal{R}}_{k}^{sa}(\Omega)\ :$ =$=\{A\in{\mathcal{F}}{\mathcal{R}}^{sa}(\Omega);\;\operatorname{index}A=k\},$
    ${\mathcal{F}}{\mathcal{R}}_{k}^{sk}(\Omega)\ :$ =$=\{A\in{\mathcal{F}}{\mathcal{R}}^{sk}(\Omega);\;\operatorname{index}A=k\},$
    ${\mathcal{F}}{\mathcal{R}}^{sa}(X)\ :$ $={\mathcal{F}}{\mathcal{R}}^{sa}(\Omega_{X}),\quad{\mathcal{F}}{\mathcal{R}}_{k}^{sa}(X)\ :={\mathcal{F}}{\mathcal{R}}_{k}^{sa}(\Omega_{X}),$
    ${\mathcal{F}}{\mathcal{R}}^{sk}(X)\ :$ $={\mathcal{F}}{\mathcal{R}}^{sk}(\Omega_{X}),\quad{\mathcal{F}}{\mathcal{R}}_{k}^{sk}(X)\ :={\mathcal{F}}{\mathcal{R}}_{k}^{sk}(\Omega_{X}).$
  • (c)
    Let $A\colon X\leadsto Y$ be a $h$-symmetric linear relation with $h\neq 0$. Define $Q_{A}\colon\operatorname{dom}(A)\times\operatorname{dom}(A)\to{\mathbb{K}}$ by $Q_{A}(x,y)\ :=\Omega(x,Ay)$ for all $x,y\in\operatorname{dom}(A)$. Then $Q_{A}$ is a well-defined sesquilinear form. We call $Q_{A}$ the associated form of $A$.
Example 2.5.
(a) We have ${\mathcal{R}}^{sk}({\mathbb{R}})=\{{\mathbb{R}}\times\{0\},\{0\}\times{\mathbb{R}}\}$.
(b) We have
${\mathcal{R}}^{sa}(\Omega)\cap{\mathcal{R}}^{sk}(\Omega)\supset\{X\times\{0\},\{0\}\times Y\}.$
The following lemma characterize the linear relations which are both symmetric and skew symmetric.
Lemma 2.6.
Let $X$ and $Y$ be two vector spaces over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{K}}$ be a nondegenerate sesquilinear form and $h\in{\mathbb{K}}$ be a number. Let $A\colon X\leadsto Y$ be a linear relation. Then the folllowing hold.
  • (a)
    $A$ is symmetric and skew symmetric with respect to $\Omega$ if and only if $A^{\Omega}\supset\operatorname{dom}(A)\times\operatorname{im}A$, if and only if $\operatorname{dom}(A)\subset(\operatorname{im}A)^{\Omega,l}$.
  • (b)
    Assume that $A$ is symmetric and skew symmetric with respect to $\Omega$. Then $A$ is a maximal (skew) symmetric linear relation with respect to $\Omega$ if and only if $A=\operatorname{dom}(A)\times\operatorname{im}A$, $\operatorname{dom}(A)=(\operatorname{im}A)^{\Omega,l}$ and $\operatorname{im}A=(\operatorname{dom}(A))^{\Omega,r}$, if and only if $A$ is selfadjoint and skew-adjoint with respect to $\Omega$.
Proof.
(a) (i) If $A^{\Omega}\supset\operatorname{dom}(A)\times\operatorname{im}A$, we have $A\cup(-A)\subset\operatorname{dom}(A)\times\operatorname{im}A\subset A^{\Omega}$.
(ii) Assume that $A\cup(-A)\subset A^{\Omega}$ holds. Let $(x,y)$ be in $A$. Then we have $(x,y)\in A^{\Omega}$ and $(x,-y)\in A^{\Omega}$. Then we obtain $(x,0)\in A^{\Omega}$ and $(0,y)\in A^{\Omega}$. Thus we have
$\Omega(x,y)=\overline{\Omega(0,y)}=0.$
Then we have $A^{\Omega}\supset\operatorname{dom}(A)\times\operatorname{im}A$ and $\operatorname{dom}(A)\subset(\operatorname{im}A)^{\Omega,l}$.
(iii) Assume that $\operatorname{dom}(A)\subset(\operatorname{im}A)^{\Omega,l}$ holds. Since $A\cup(-A)\subset\operatorname{dom}(A)\times\operatorname{im}A$, we have $A^{\Omega}\supset\operatorname{dom}(A)\times\operatorname{im}A$. By Step a.i we have $A\cup(-A)\subset A^{\Omega}$.
(b) (i) By (a), we have $A\cup(-A)\subset\operatorname{dom}(A)\times\operatorname{im}A\subset A^{\Omega}$, and $\operatorname{dom}(A)\times\operatorname{im}A$, $\operatorname{dom}(A)\times(\operatorname{dom}(A))^{\Omega,r}$, and $(\operatorname{im}A)^{\Omega,l}\times\operatorname{im}A$ are three symmetric and skew symmetric linear relations with respect to $\Omega$ containing $A$.
(ii) If $A$ is a maximal (skew) symmetric linear relation with respect to $\Omega$, by Step b.i we have $A=\operatorname{dom}(A)\times\operatorname{im}A=\operatorname{dom}(A)\times(\operatorname{dom}(A))^{\Omega,r}=(\operatorname{im}A)^{\Omega,l}\times\operatorname{im}A$. Then we have $A=\operatorname{dom}(A)\times\operatorname{im}A$, $\operatorname{dom}(A)=(\operatorname{im}A)^{\Omega,l}$ and $\operatorname{im}A=(\operatorname{dom}(A))^{\Omega,r}$.
(iii) By definition and direct calculation we obtain the rest implications. ∎
Let $X,Y$ be two Banach spaces. Let $A\colon X\leadsto Y$ be a linear relation. By [13, (III.5.9)] we have
(20) $(-A^{*})^{-1}=A^{\perp}.$
Given a bounded nondegenerate sesquilinear form $\Omega\colon X\times Y\to{\mathbb{K}}$, we have
(21) $A^{\Omega}=(L_{\omega})^{-1}A^{\perp}=(L_{\omega})^{-1}(-A^{*})^{-1}=(R_{\Omega})^{-1}\circ A^{*}\circ L_{\Omega},$
where $L_{\omega}$ is defined by (19), $L_{\Omega}$ and $R_{\Omega}$ are defined by (1).
Let $X$ be a Banach space. We denote by $\iota_{X}\colon X\to X^{**}$ the natural embedding and $\mathrm{i}\ :=\sqrt{-1}$. Then we have $R_{\Omega}=L_{\Omega}^{*}\circ\iota_{Y}$, $L_{\Omega_{X}}=\iota_{X}$ and $R_{\Omega_{X}}=I_{X^{*}}$. By (21) we have $A^{\Omega_{X}}=A^{*}\circ\iota_{X}$. Then $A$ is skew-adjoint if and only if $\mathrm{i}A$ is selfadjoint, $A$ is $h$-symmetric if and only if $hA\subset A^{*}\circ\iota_{X}$, and $A$ is $h$-selfadjoint if and only if $hA=A^{*}\circ\iota_{X}$.
The number $|h|$ can only takes value of $1$ in the non-trivial case. More precisely, we have the following result.
Lemma 2.7.
[…]
Proof.
Since $A^{\Omega}=hA$, we have
$A\subset A^{\Omega\Omega}=(hA)^{\Omega}=\bar{h}A^{\Omega}=|h|^{2}A.$
[…]
A symmetric linear relation has the following property.
Lemma 2.8.
Let $X$ and $Y$ be two vector space over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{K}}$ be a nondegenerate sesquilinear form with respect to $\Omega$. Let $A\colon X\leadsto Y$ be a symmetric linear relation. Then we have $\Omega(x,y)=0$ for $(x,y)\in(\ker A\times\operatorname{im}A)\cup(\operatorname{dom}(A)\times A0)$.
Proof.
Since $(x,y)\in A$ and $A\subset A^{\Omega}$, we have $(x,y)\in A^{\Omega}$. Since $x=0$ or $y=0$ holds, by (14) we have $\Omega(x,y)=0$. ∎
For a maximal symmetric linear relation, we have the following.
Proposition 2.9.
Let $X$ and $Y$ be two vector space over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{K}}$ be a nondegenerate sesquilinear form. Let $A\colon X\leadsto Y$ be a maximal symmetric linear relation. Then we have
(22) $\ker A=\ker A^{\Omega}=(\operatorname{im}A)^{\Omega,l},$ $\quad A0=A^{\Omega}0=(\operatorname{dom}(A))^{\Omega,r},$
(23) $\operatorname{im}A^{\Omega}\subset((\operatorname{im}A)^{\Omega,l})^{\Omega,r},$ $\quad\operatorname{dom}A^{\Omega}\subset((\operatorname{dom}A)^{\Omega,r})^{\Omega,l}.$
Proof.
Let $\omega$ be defined by (14).
1. Let $x\in(\operatorname{im}A)^{\Omega,l}$. By (14) we have $(x,0)\in A^{\omega}$. Note that we have $\omega((x,0),(x,0))=0$. Then the subspace $A+\operatorname{span}\{(x,0)\}$ of $X\times Y$ is isotropic. Since $A$ is maximal isotropic, we have $(x,0)\in A$ and $x\in\ker A$. Then we have $\ker A\supset(\operatorname{im}A)^{\Omega,l}$. By Lemma 2.8 we have $\ker A\subset(\operatorname{im}A)^{\Omega,l}$ and then $\ker A=(\operatorname{im}A)^{\Omega,l}$. Apply the result for $A^{-1}$, we have $A0=(\operatorname{dom}(A))^{\Omega,r}$.
2. Let $(x,y)\in A^{\Omega}$. Then we have
$0=\omega((x,y),(0,y_{1}))=\Omega(x,y_{1})$
for all $y_{1}\in A0$. By Step 1 we have $x\in(A0)^{\Omega,l}=((\operatorname{dom}(A))^{\Omega,r})^{\Omega,l}.$ Then we obtain $\operatorname{dom}(A^{\Omega})\subset((\operatorname{dom}(A))^{\Omega,r})^{\Omega,l}$. Apply the result for $A^{-1}$, we have $\operatorname{im}A^{\Omega}\subset((\operatorname{im}A)^{\Omega,l})^{\Omega,r}$.
3. Let $(x,y)\in A^{\Omega}$ and $x\in\operatorname{dom}(A)$. Then there is a $y_{2}\in Y$ such that $(x,y_{2})\in A$. Then there holds $(0,y-y_{2})=(x,y_{2})-(x,y)\in A^{\Omega}$. For each $(x_{3},y_{3})\in A$ we have
$0=\omega((x_{3},y_{3}),(0,y-y_{2}))=\Omega(x_{3},y-y_{2}).$
By Step 1 we have $y-y_{2}\in(\operatorname{dom}(A))^{\Omega,r}=A0$ and $y\in y_{2}+A0=Ax$. Then we get $A^{\Omega}\cap(\operatorname{dom}(A)\times Y)=A$ and $A^{\Omega}0=A0$. Apply the result for $A^{-1}$, we have $\ker A^{\Omega}=\ker A$. ∎
The index of a ±$\pm 1$ selfadjoint linear relation is always non-positive. More precisely, we have the following results.
Lemma 2.10.
Let $X$ and $Y$ be two vector space over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{C}}$ be a nondegenerate sesquilinear form. Let $A\colon X\leadsto Y$ be an ±$\pm 1$ symmetric linear relation with respect to $\Omega$. Then the following hold.
(a) We have $\operatorname{index}A\leq 0$ and $\dim\ker A<+\infty$ if one of $\dim\ker A$ and $\dim Y/\operatorname{im}A$ is finite.
(b) The equality $\operatorname{index}A=0$ holds if and only if $\dim\ker A=\dim\ker A^{\Omega}$ and $\operatorname{im}A=((\operatorname{im}A)^{\Omega,l})^{\Omega,r}$.
(c) The relation $A$ is ±$\pm 1$ selfadjoint if the equality $\operatorname{index}A=0$ holds.
Proof.
Let $\omega$ be defined by (14).
(a) Note that (±$(\pm A)+X\times\{0\}=X\times\operatorname{im}A$. If one of $\dim\ker A$ and $\dim Y/\operatorname{im}A$ is finite, by [5, Lemma 1.2.8] we have
$\dim\ker A$ $=\dim(A\cap(X\times\{0\}))$
$\leq\dim((\pm A)^{\omega}\cap(X\times\{0\})^{\omega})$
$\leq\dim(X\times Y)/((\pm A)+X\times\{0\})$
$=\dim Y/\operatorname{im}A.$
Then we have $\operatorname{index}A\leq 0$ and $\dim\ker A<+\infty$.
(b) By [5, Lemma 1.2.8], $\operatorname{index}A=0$ holds if and only if $\dim\ker A=\dim\ker A^{\omega}$ and $(X\times\operatorname{im}A)^{\omega\omega}=X\times\operatorname{im}A$, if and only if $\dim\ker A=\dim\ker A^{\omega}$ and $\operatorname{im}A=(\operatorname{im}A^{\Omega,l})^{\Omega,r}$.
(c) For ${\mathbb{K}}={\mathbb{C}}$, we note that ±$\sqrt{\pm 1}A$ is symmetric and $\operatorname{index}(\sqrt{\pm 1}A)=0$. By [4, Proposition 1 ], ±$\sqrt{\pm 1}A$ is selfadjoint. Then $A$ is ±$\pm 1$ selfadjoint.
For ${\mathbb{K}}={\mathbb{R}}$, by the complex case the complexified linear rellation ±$\sqrt{\pm 1}A\otimes_{{\mathbb{R}}}{\mathbb{C}}$ is selfadjoint. Then $A$ is ±$\pm 1$ selfadjoint. ∎
We have the following generalization of [7, Lemma 2.2], where the full proof of the corresponding result in N.J. Kalton, R. C. Swanson [12] was given.
Proposition 2.11.
Let $X$ be a Banach space. Let $A\colon X\leadsto X^{*}$ be a ±$\pm 1$ selfadjoint linear relation. Then the following holds.
  • (a)
    Assume that $A$ is semi-Fredholm. Then $\operatorname{index}A=0$ holds if and only if $\operatorname{im}\iota_{X}\supset\operatorname{dom}(A^{*})$, if and only if $\operatorname{im}A=\operatorname{im}A^{*}$.
  • (b)
    We have $A^{*}0=A0$, and $\iota_{X}$ induces a linear isomorphism $X/\operatorname{dom}(A)\to\operatorname{im}\iota/(\operatorname{im}\iota\cap\operatorname{dom}(A^{*}))$.
Proof.
1. Since ±$\pm A=A^{\omega}=A^{*}\circ\iota_{X}$ and $\iota_{X}$ is an injection, by [15, Exercise B.2.11], there are exact sequences
(24) $\begin{split}\{0\}&\to\ker A\to\ker A^{*}\to\operatorname{dom}(A^{*})/(\operatorname{dom}(A^{*})\cap\operatorname{im}\iota_{X})\\ &\to X^{*}/\operatorname{im}A\to X^{*}/\operatorname{im}A^{*}\to\{0\},\end{split}$
(25) $\begin{split}\{0\}&\to A^{*}0\to\ker A0\to\{0\}\\ &\to\operatorname{im}\iota/(\operatorname{im}\iota\cap\operatorname{dom}(A^{*}))\to X/\operatorname{dom}(A)\to\{0\},\end{split}$
2. Proof of (a).
Since $A$ is semi-Fredholm, by Lemma 2.10 and the proof of [13, Theorem IV.5.13], we have
$\operatorname{index}A$ $=\dim\ker A-\dim\ker A^{*}$
$=\dim X^{*}/\operatorname{im}A^{*}-\dim X^{*}/\operatorname{im}A\leq 0.$
By (24), $\operatorname{index}A=0$ holds if and only if the injection $\ker A\to\ker A^{*}$ is an isomorphism, if and only if $\operatorname{dom}(A^{*})\cap\operatorname{im}\iota_{X}=\operatorname{dom}(A^{*})$, if and only if $\operatorname{im}\iota_{X}\supset\operatorname{dom}(A^{*})$, if and only if the surjection $X^{*}/\operatorname{im}A\to X^{*}/\operatorname{im}A^{*}$ is an isomorphism, if and only if $\operatorname{im}A=\operatorname{im}A^{*}$.
3. By (24) we obtain (b). ∎
For symplectic linear subspace, we have the following.
Lemma 2.12.
Let $(X,\omega)$ be a symplectic vector space with a symplectic linear subspace $X_{0}$. Then $\operatorname{im}J(\cdot)|_{X_{0}}\subset\operatorname{im}J_{\omega|_{X_{0}\times X_{0}}}$ holds if and only if
$X=X_{0}\oplus X_{0}^{\omega}.$
In this case we have $X_{0}^{\omega\omega}=X_{0}$.
Proof.
Since $X_{0}$ is symplectic, $J_{\omega|_{X_{0}\times X_{0}}}$ is injective and we have $X_{0}\cap X_{0}^{\omega}=\{0\}$. Then we have
$\operatorname{im}J(\cdot)|_{X_{0}}\subset\operatorname{im}J_{\omega|_{X_{0}\times X_{0}}}$
$\Leftrightarrow$ $\forall x\in X,\exists x_{0}\in X_{0}\text{ such that }J(x)|_{X_{0}}=J_{\omega|_{X_{0}\times X_{0}}}(x_{0})$
$\Leftrightarrow$ $\forall x\in X,\exists x_{0}\in X_{0}\text{ such that }x-x_{0}\in X_{0}^{\omega}$
$\Leftrightarrow$ $X\subset X_{0}+X_{0}^{\omega}$
$\Leftrightarrow$ $X=X_{0}\oplus X_{0}^{\omega}.$
In this case, we have $X_{0}^{\omega\omega}\cap X_{0}^{\omega}=\{0\}$. By [5, Lemma A.1.1], we have
$X_{0}^{\omega\omega}$ $=X_{0}^{\omega\omega}\cap(X_{0}+X_{0}^{\omega})$
$=X_{0}+X_{0}^{\omega\omega}\cap X_{0}^{\omega}=X_{0}.$
Corollary 2.13.
(cf. [7, Lemma 2.7]) Let $(X,\omega)$ be a symplectic Banach space with a linear subspace $X_{0}$. Then the following hold.
(a) If $(X_{0},\omega|_{X_{0}\times X_{0}})$ is a strong symplectic Banach space, we have
$X=X_{0}\oplus X_{0}^{\omega}.$
(b) If $X=X_{0}\oplus X_{0}^{\omega}$ holds and $(X,\omega)$ is a strong symplectic Banach space, $(X_{0},\omega|_{X_{0}\times X_{0}})$ is a strong symplectic Banach space.
Proof.
(a) Since $(X_{0},\omega|_{X_{0}\times X_{0}})$ is a strong symplectic Banach space, we have $\operatorname{im}J_{\omega|_{X_{0}\times X_{0}}}=X_{0}^{*}\supset\operatorname{im}J(\cdot)|_{X_{0}}$. By Lemma 2.12, we have $X=X_{0}\oplus X_{0}^{\omega}$.
(b) Since $(X,\omega)$ is a strong symplectic Banach space, we have $\operatorname{im}J=X^{*}$. By Hahn-Banach theorem, we have $X_{0}^{*}\subset\operatorname{im}J(\cdot)|_{X_{0}}$. Since $X=X_{0}\oplus X_{0}^{\omega}$ holds, by Lemma 2.12, we have $\operatorname{im}J(\cdot)|_{X_{0}}\subset\operatorname{im}J_{\omega|_{X_{0}\times X_{0}}}$. Since $\operatorname{im}J(\cdot)|_{X_{0}}$ and $\operatorname{im}J_{\omega|_{X_{0}\times X_{0}}}$ are subsets of $X_{0}^{*}$, we have $X_{0}^{*}=\operatorname{im}J(\cdot)|_{X_{0}}=\operatorname{im}J_{\omega|_{X_{0}\times X_{0}}}$, i.e. $X_{0}$ is a strong symplectic Banach space. ∎
2.3. Perturbed Morse index
Before study the stability theorems of maximal isotropic spaces, we study the perturbed Morse index for symmetric forms with varying domains in a fixed Banach space. Note that in general one can not make the domains fixed.
Let $V$ be a vector space over ${\mathbb{K}}$ and $Q\colon V\times V\to{\mathbb{K}}$ be a symmetric form. For each subset $\lambda$ of $V$, we denote by
$\lambda^{Q}\ :=\lambda^{Q,l}=\lambda^{Q,r}.$
Then $m^{\pm}(Q)$ and $m^{0}(Q)$ denotes the Morse positive (or negative) index and the nullity of $Q$ respectively.
Definition 2.14.
(a) Let $X$ be a Banach space over ${\mathbb{K}}$ with a closed linear subspace $V$. Let $Q\colon V\times V\to{\mathbb{K}}$ be a symmetric form. If
(26) $\|Q\|\ :=\begin{cases}\sup\limits_{x,y\in V\setminus\{0\}}\frac{|Q(x,y)|}{\|x\|\|y\|}<+\infty&\text{ if }V\neq\{0\},\\ 0&\text{ if }V=\{0\},\end{cases}$
we call $(Q,V)$ a bounded symmetric pair of $X$.
(b) Let $X$ be a Banach space over ${\mathbb{K}}$ with a closed linear subspace $V$. Let $Q\colon V\times V\to{\mathbb{K}}$ be a bounded semi-definite symmetric form. The reduced minimum modulus $\gamma(Q)$ of $Q$ is defined by
(27) $\gamma(Q)\ :=\begin{cases}\inf\limits_{x\in V\setminus V^{Q}}\frac{|Q(x,x)|}{(\operatorname{dist}(x,V^{Q}))^{2}}&\text{ if }V\neq V^{Q},\\ 0&\text{ if }V=V^{Q},\end{cases}$
(c) Let $X$ be a Banach space with two symmetric pairs $(Q,V)$ and $(R,W)$. Let $c\geq 0$ be a real number. We define the $c$-gap $\delta_{c}(Q,R)$ between $R$ and $Q$ to be the infimum of the non-negative number $\delta$ such that
(28) $\begin{split}|Q(x,y)-&R(u,v)|\leq\delta(\|u\|+\|x\|)(\|v\|+\|y\|)+\\ &c\left((\|u\|+\|x\|)\|v-y\|+\|u-x\|(\|v\|+\|y\|)\right)\end{split}$
for all $x,y\in V$ and $u,v\in W$.
We have the following properties of the $c$-gap.
Lemma 2.15.
Let $X$ be a Banach space with two symmetric pairs $(Q,V)$ and $(R,W)$. Let $c\geq 0$ be a real number. Then the following hold.
  • (a)
    The equation (28) holds for $\delta=\delta_{c}(Q,R)$ and $\delta_{c}(Q,R)=\delta_{c}(R,Q)$.
  • (b)
    Assume that $\delta_{c}(Q,R)$ is finite. Then $(Q,V)$ and $(R,W)$ are bounded and we have
    (29) $\max\{\|Q\|,\|R\|\}\leq\delta_{c}(Q,R)+2c.$
  • (c)
    Assume that $(Q,V)$ and $(R,W)$ are bounded. Then we have
    (30) $0\leq\delta_{c}(Q,R)\leq\max\{\|Q\|,\|R\|\}.$
  • (d)
    The function $\delta_{c}(Q,R)$ is decreasing on $c\geq 0$.
  • (e)
    Let $M$ and $N$ be linear subspace of $V$ and $W$ respectively. Then we have
    (31) $\delta_{c}(Q|_{M},R|_{N})\leq\delta_{c}(Q,R).$
  • (f)
    Assume that $(Q,V)$ and $(R,W)$ are bounded. If $V=W$ holds, we have
    (32) $\frac{\|R-Q\|}{4}\leq\delta_{c}(Q,R)\leq\|R-Q\|$
    for $c\geq\|Q\|$.
Proof.
(a) By definition, we have $\delta_{c}(Q,R)=\delta_{c}(R,Q)$.
Let $\varepsilon>0$ be arbitrary. Eor $x,y\in V$ and $u,v\in W$, (28) holds for $\delta=\delta_{c}(Q,R)+\varepsilon$. On letting $\varepsilon\to 0$, we have (28) holds for $x,y\in V$, $u,v\in W$ and $\delta=\delta_{c}(Q,R)$.
(b) Take $u=v=0$ in (28), we have
$|Q(x,y)|\leq(\delta_{c}(Q,R)+2c)\|x\|\|y\|$
for all $x,y\in V$. By definition, we have $\|Q\|\leq\delta_{c}(Q,R)+2c$. Change the role of $Q$ and $R$ we have $\|R\|\leq\delta_{c}(Q,R)+2c$.
(c) By definition, we have $\delta_{c}(Q,R)=\delta_{c}(R,Q)\geq 0$. Let $x,y\in V$ and $u,v\in W$ be elements in $X$. Then we have
$|Q(x,y)$ $-R(u,v)|\leq|Q(x,y)|+|R(u,v)|$
$\leq\|Q\|\|x\|\|y\|+\|R\|\|u\|v\|$
$\leq\max\{\|Q\|,\|R\|\}(\|u\|+\|x\|)(\|v\|+\|y\|).$
By definition we obtain (30).
(d) Let $c$, $d$ be non-negative numbers with $c\leq d$. Then we have
$|Q(x,y)-$ $R(u,v)|\leq(\delta_{c}(P,Q))(\|u\|+\|x\|)(\|v\|+\|y\|)+$
$d\left((\|u\|+\|x\|)\|v-y\|+\|u-x\|(\|v\|+\|y\|)\right)$
for all $x,y\in V$ and $u,v\in W$. By definition, we have $\delta_{d}(P,Q)\leq\delta_{c}(P,Q)$.
(e) Since (28) holds for all $x,y\in M$ and $u,v\in N$, by definition, we have $\delta_{c}(Q|_{M},R|_{N})\leq\delta_{c}(P,Q)$.
(f) Let $x$, $y$, $u$, $v$ be in $V=W$. Since $c\geq\|Q\|$ holds, we have
$|Q(x,y)-$ $R(u,v)|\leq|Q(u,v)-R(u,v)|+|Q(x,y-v)|+|Q(x-u,v)|$
$\leq$ $\|Q-R\|\|u\|\|v\|+\|Q\|(\|u\|\|v-y\|+\|u-x\|\|v\|)$
$\leq\|Q-R\|(\|u\|+\|x\|)(\|v\|+\|y\|)+$
$c\left((\|u\|+\|x\|)\|v-y\|+\|u-x\|(\|v\|+\|y\|)\right).$
By definition, we have
$\delta_{c}(Q,R)\leq\|R-Q\|.$
By (28), we have
$|Q(x,y)-R(x,y)|\leq 4(\delta_{c}(Q,R))\|x\|\|y\|$
for all $x,y\in V=W$. Then we have
$\delta_{c}(Q,R)\geq\frac{\|R-Q\|}{4}.$
The main result of this subsection is the following.
Proposition 2.16.
Let $X$ be a Banach space with two bounded symmetric pairs $(Q,V)$ and $(R,W)$. Let $c\geq 0$ be a real number. Let $\alpha$ be a linear subspace of $V$ with $\dim\alpha=k\in[1,+\infty)$, $h$ be in $\{1,-1\}$ and $hQ|_{\alpha}$ be positive definite. Assume that
(33) $k(\delta_{c}(Q,R)(2+\delta(V,W))+2c\delta(V,W))(2+\delta(V,W))<\gamma(Q|_{\alpha}).$
Then we have
(34) $m^{+}(hR)\geq\dim\alpha.$
Proof.
Take an orthogonal bases $v_{1},\ldots,v_{k}$ of $\alpha$ with respect to $hQ$ such that $\|v_{i}\|=1$ for each $i=1,\ldots,k$. For each $\varepsilon>0$ and $i=1,\ldots,k$, there is a $w_{i}\in W$ with $\|v_{i}-w_{i}\|\leq(1+\varepsilon)\delta(V,W)$.
By (28), we have
$|R(w_{i},w_{j})$ $-Q(v_{i},v_{j})|\leq\delta_{c}(Q,R)(\|w_{i}\|+\|v_{i}\|)(\|w_{j}\|+\|v_{j}\|)+$
$c\left(\|v_{i}-w_{i}\|(\|v_{j}\|+\|w_{j}\|)+\|v_{j}-w_{j}\|(\|v_{i}\|+\|w_{i}\|)\right)$
$\leq$ $\delta_{c}(Q,R)(2+(1+\varepsilon)\delta(V,W))^{2}+$
$2c(1+\varepsilon)\delta(V,W)(2+(1+\varepsilon)\delta(V,W)):\ =C(\varepsilon).$
Set $A\ :=(Q(v_{i},v_{j}))_{1\leq i,j\leq k}$ and $B\ :=(R(w_{i},w_{j}))_{1\leq i,j\leq k}$. Then we have $hA\geq\gamma(Q|_{\alpha})I_{k}>0$. Since $kC(0)<\gamma(Q|_{\alpha})$, for $\varepsilon>0$ sufficiently small, we have
$h(B-A)\geq-kC(\varepsilon)I_{k}>-\gamma(Q|_{\alpha})I_{k}.$
Since $hA\geq\gamma(Q|_{\alpha})I_{k}>0$, we have $hB>0$. Then $\{w_{1},\ldots,w_{k}\}$ is a linearly independent subset and we have $m^{+}(hR)\geq k$. ∎
2.4. Stability theorems of maximal isotropic subspaces
In this subsection we study the stability theorems of maximal isotropic subspaces. Firstly we recall the concept of symplectic reduction.
Let $(X,\omega)$ be a symplectic vector space with an isotropic subspace $\lambda$. Then the induced skew symmetric form on the vector space $\lambda^{\omega}/\lambda$ is defined by
(35) $\widetilde{\omega}(x+\lambda,y+\lambda)\ :=\ \omega(x,y)\text{ for all }x,y\in\lambda^{\omega}.$
The space $(\lambda^{\omega}/\lambda,\widetilde{\omega})$ is symplectic if and only if $\lambda=\lambda^{\omega\omega}$. We call $(\lambda^{\omega}/\lambda,\widetilde{\omega})$ the symplectic reduction of $X$ by $\lambda$ and denote by $\pi_{\lambda}\colon\lambda^{\omega}\to\lambda^{\omega}/\lambda$ the canonical projection. Let $\alpha$ be a linear subspace of $X$. The symplectic reduction of $\alpha$ by $\lambda$ is defined by
(36) $\pi_{\lambda}(\alpha)\ :=\ \bigl{(}(\alpha+\lambda)\cap\lambda^{\omega}\bigr{)}/\lambda\ =\ \bigl{(}\alpha\cap\lambda^{\omega}+\lambda\bigr{)}/\lambda.$
We need the following properties of maximal isotropic subspaces.
Proposition 2.17.
Let $(X,\omega)$ be a symplectic vector space with an isotropic subspace $\lambda$. Then the following hold.
(a) If $\lambda$ is a maximal isotropic subspace of $(X,\omega)$, we have $\lambda=\lambda^{\omega\omega}$.
(b) $\lambda$ is a maximal isotropic subspace of $(X,\omega)$ if and only if
(37) $\lambda=\{x\in\lambda^{\omega};\;\omega(x,x)=0\}.$
(c) Assume that $(X,\omega)$ is a complex symplectic vector space. Then $\lambda$ is a maximal isotropic subspace of $(X,\omega)$ if and only if $\mathrm{i}\tilde{\omega}$ is definite on $\lambda^{\omega}/\lambda$.
(d) Assume that $(X,\omega)$ is a real symplectic vector space. Then $\lambda$ is a maximal isotropic subspace of $(X,\omega)$ if and only if $\lambda=\lambda^{\omega}$.
Proof.
(a) Since $\lambda\subset\lambda^{\omega}$ holds, we have
$\lambda\subset\lambda^{\omega\omega}\subset\lambda^{\omega}=\lambda^{\omega\omega\omega}.$
Since $\lambda$ is a maximal isotropic subspace of $(X,\omega)$, we have $\lambda=\lambda^{\omega\omega}$.
(b) (i) Necessity.
Since $\lambda$ is isotropic, we have $\lambda\subset\{x\in\lambda^{\omega};\;\omega(x,x)=0\}$. Let $x\in\lambda^{\omega}$ be such that $\omega(x,x)=0$. Assume that $x\notin\lambda$. Since $\omega(x,x)=0$, the subspace $\lambda\oplus\operatorname{span}\{x\}\supsetneq\lambda$ is an istropic subspace of $(X,\omega)$. This contradicts to the fact that $\lambda$ is a maximal isotropic subspace of $(X,\omega)$. Then we have $\lambda\supset\{x\in\lambda^{\omega};\;\omega(x,x)=0\}$. Thus we obtain (37).
(ii) Sufficiency.
Let $\mu\supset\lambda$ be isotropic. Then we have $\mu\subset\mu^{\omega}\subset\lambda^{\omega}$. By (37), we have $\mu\subset\{x\in\lambda^{\omega};\;\omega(x,x)=0\}=\lambda$. Then we have $\lambda=\mu$. Since $\mu$ is arbitrary, $\lambda$ is maximal isotropic.
(c) (i) Necessity.
By (a), $\tilde{\omega}$ is a symplectic form on $\lambda^{\omega}/\lambda$. If $\mathrm{i}\tilde{\omega}$ is not definite on $\lambda^{\omega}/\lambda$, by (b), there are $x,y\in\lambda^{\omega}$ such that $\mathrm{i}\omega(x,x)>0$ and $\mathrm{i}\omega(y,y)<0$. Then there is an $a\in{\mathbb{C}}\setminus\{0\}$ such that $\mathrm{i}\omega(ax+y,ax+y)=0$. By (b), we have $ax+y\in\lambda$. Then we have
$\omega(y,y)=\omega(ax+y-ax,ax+y-ax)=|a|^{2}\omega(x,x).$
From this we obtain that $|a|^{2}\mathrm{i}\omega(x,x)=\mathrm{i}\omega(y,y)$, which contridicts to the facts $\mathrm{i}\omega(x,x)>0$ and $\mathrm{i}\omega(y,y)<0$.
(ii) Sufficiency.
Since $\mathrm{i}\tilde{\omega}$ is definite on $\lambda^{\omega}/\lambda$, (37) holds. By (b), $\lambda$ is maximal isotropic.
(d) (i) Necessity.
Since $X$ is real, we have $\omega(x,x)=0$ for all $x\in X$. By (b), $\lambda$ is a Lagrangian subspace of $(X,\omega)$.
(ii) Sufficiency.
Since $\lambda$ is a Lagrangian subspace of $(X,\omega)$, it is maximal isotropic. ∎
Based on Proposition 2.17, we can define the sign of a maximal isotropic subspace.
Definition 2.18.
Let $(X,\omega)$ be a symplectic vector space and $\lambda$ be a maximal isotropic subspace of $(X,\omega_{0})$. We define the sign $h_{\lambda}$ by
(38) $h_{\lambda}\ :=\begin{cases}\pm 1&\text{ if }\pm\mathrm{i}\omega|_{\lambda}\text{ is positive definite},\\ 0&\text{ if }\lambda=\lambda^{\omega}.\end{cases}$
and the reduced minimum modulus $\gamma_{\lambda}$ of $\lambda$ by
(39) $\gamma_{\lambda}\ :=\begin{cases}\inf\limits_{x\in\lambda^{\omega}\setminus\lambda}\frac{\mathrm{i}h_{\lambda}\omega(x,x)}{(\operatorname{dist}(x,\lambda))^{2}}&\text{ if }\lambda\subsetneq\lambda^{\omega}\\ 0&\text{ if }\lambda=\lambda^{\omega}\end{cases}$
respectively.
By definition, a maximal isotropic subspace $\lambda$ is Lagrangian if and only if $h_{\lambda}=0$.
The starting point of our stability theorems of maximal isotropic subspaces is the following result of P. Hess and T. Kato [11, Lemma 1]).
Lemma 2.19.
Let $X$ be a Banach space with three closed linear subspaces $M$, $N$ and $N^{\prime}$, where $N^{\prime}\subset N$. If there hold that
$(1+\delta(N,M))(1+\delta(M,N^{\prime}))<2,$
we have $N^{\prime}=N$.
Proof of Theorem 1.1.
Assume that $\mu$ is not a a maximal isotropic subspace of $(X,\omega)$ with $h_{\mu}\in\{h_{\lambda},0\}$. Then we have $\mu\subsetneq\mu^{\omega}$. By Proposition 2.17.b, there is an $x_{1}\in\mu^{\omega}\setminus\mu$ such that
$\mathrm{i}h_{\lambda}\omega(x_{1},x_{1})\leq 0.$
Since $\mu$ is closed, by [13, Lemma III.1.12], for each $\varepsilon\in(0,1)$, there is a $y_{1}\in\mu$ such that
$\operatorname{dist}(x_{1},\mu)>(1-\varepsilon)\|x_{1}-y_{1}\|>0.$
Set $x\ :=\frac{x_{1}-y_{1}}{\|x_{1}-y_{1}\|}$. Then we have $x\in\mu^{\omega}\setminus\mu$, $\|x\|=1$, $\mathrm{i}h_{\lambda}\omega(x,x)\leq 0$, and
$\operatorname{dist}(x,\mu)=\operatorname{dist}(\frac{x_{1}}{\|x_{1}-y_{1}\|},\mu)>1-\varepsilon>0.$
By definition, there is a $y\in\lambda^{\omega_{0}}$ such that
$\|x-y\|<\delta(\mu^{\omega},\lambda^{\omega_{0}})+\varepsilon.$
By Lemma 2.21, we have
$\operatorname{dist}(y,\lambda)\leq$ $\gamma_{\lambda}^{-\frac{1}{2}}(\mathrm{i}h_{\lambda}\omega_{0}(y,y))^{\frac{1}{2}}$
$=$ $\left(\mathrm{i}h_{\lambda}\gamma_{\lambda}^{-1}\left(\omega(x,x)+\omega_{0}(x,x)-\omega(x,x)+\right.\right.$
$\left.\left.\omega_{0}(y-x,y)+\omega_{0}(x,y-x)\right)\right)^{\frac{1}{2}}$
$\leq$ $\left(|h_{\lambda}|\gamma_{\lambda}^{-1}\left(\|L_{\omega}-L_{\omega_{0}}\|+\|L_{\omega_{0}}\|\|x-y\|(2+\|x-y\|)\right)\right)^{\frac{1}{2}}.$
By [13, Lemma IV.2.2], we have
$(1+\delta(\lambda,\mu))$ $(\|x-y\|+\operatorname{dist}(y,\lambda))\geq(1+\delta(\lambda,\mu))\operatorname{dist}(x,\lambda)$
$\geq\operatorname{dist}(x,\mu)-\delta(\lambda,\mu)$
$>1-\varepsilon-\delta(\lambda,\mu).$
Then we have
$\delta(\mu^{\omega},\lambda^{\omega_{0}})+$ $\varepsilon+\left(|h_{\lambda}|\gamma_{\lambda}^{-1}\left(\|L_{\omega}-L_{\omega_{0}}\|+\|L_{\omega_{0}}\|(\delta(\mu^{\omega},\lambda^{\omega_{0}})+\varepsilon)\right.\right.$
$\left.\left.(2+\delta(\mu^{\omega},\lambda^{\omega_{0}})+\varepsilon)\right)\right)^{\frac{1}{2}}>\frac{1-\varepsilon-\delta(\lambda,\mu)}{1+\delta(\lambda,\mu)}.$
On letting $\varepsilon\to 0$ we obtain
(40) $\delta(\mu^{\omega},\lambda^{\omega_{0}})+l(\mu,\omega)\geq\frac{1-\delta(\lambda,\mu)}{1+\delta(\lambda,\mu)}.$
This contradicts to (2). So $\mu$ is a maximal isotropic subspace of $(X,\omega)$ with $h_{\mu}\in\{h_{\lambda},0\}$. ∎
Proof of Theorem 1.2.
1. Denote by
$K_{1}\ :=$$\{s\in J;\;\lambda(s)=\lambda^{\omega(s)\omega(s)}\},$
$K_{2}\ :=$$\{s\in J;\;\lambda(s)=\lambda^{\omega(s)}\}.$
Then $K_{1}$ and $K_{2}$ are closed subsets of $J$. By Proposition 2.17.a, we have $s_{0}\in K_{1}$. By Lemma 2.19, $K_{1}$ and $K_{2}$ are open subsets of $J$. Since $J$ is connected, we have $K_{1}=J$ and $K_{2}$ is either $J$ or empty set. Then we have $K_{2}=J$ if $s_{0}\in K_{2}$, and in this case we have $h_{\lambda(s)}=h_{\lambda(s_{0})}=0$ holds for each $s\in J$.
2. Assume that $K_{2}=\emptyset$ holds. Then we have $h_{\lambda(s_{0})}\neq 0$. By Proposition 2.17.d, we have ${\mathbb{K}}={\mathbb{C}}$. Denote by
$K_{3}\ :=\{s\in J;\;m^{-}(\mathrm{i}h_{\lambda(s_{0})}\omega_{s}|_{\lambda(s)^{\omega}(s)})=0\}.$
By Proposition 2.16, $K_{3}$ is closed. Since $K_{1}=J$, by Proposition 2.17.c, we have
$K_{3}=$$\{s\in J;\;\lambda(s)\text{ is a maximal isotropic subspace of }$
$(X,\omega(s))\text{ and }h_{\lambda(s)}\in\{h_{\lambda(s_{0})},0\}\}$
$=$ $\{s\in J;\;\lambda(s)\text{ is a maximal isotropic subspace of }$
$(X,\omega(s))\text{ and }h_{\lambda(s)}=h_{\lambda(s_{0})}\}.$
Then we have $s_{0}\in K_{3}$. By Theorem 1.1, $K_{3}$ is open. Since $K_{3}\subset J$ is a nonempty set and $J$ is connected, we have $K_{3}=J$. ∎
2.5. Stability theorems in strong symplectic Banach spaces
In strong symplectic Banach spaces, some conditions in stability theorems of maximal isotropic subspaces hold. Then these stability theorems can be simplified.
If $(X,\omega(s))$ is a strong symplectic Banach space for each $s\in J$, the continuity of the family $\{\lambda(s)^{\omega(s)}\}_{s\in J}$ follows from that of the family $\{\lambda(s)\}_{s\in J}$ and the following estimate (cf. [13, Theorem IV.2.29] for a similar result).
Lemma 2.20.
Let $X$, $Y$ be two Banach spaces. Let $A\in{\mathcal{B}}(X,Y)$ be a bounded operator with bounded inverse. Assume that $B\in{\mathcal{B}}(X,Y)$ satisfies $\|A-B\|\|A^{-1}\|<1$. Let $M$, $N$ be two closed linear subspaces of $X$. Then the spaces $AM$ and $BN$ are closed, and we have
(41) $\delta(AM,BN)\leq\|A^{-1}\|(\|A-B\|+\|B\|\delta(M,N)).$
Proof.
(a) Since $A,A^{-1},B\in{\mathcal{B}}(X,Y)$ and $\|A-B\|\|A^{-1}\|<1$, we have $B^{-1}\in{\mathcal{B}}(X,Y)$. Since $M,N$ are closed, $AM$ and $BN$ are closed.
If $M=\{0\}$, we have $\delta(AM,BN)=0$ and (41) holds.
Now we assume $M\neq\{0\}$. For each $x\in M\setminus\{0\}$ we have
$\operatorname{dist}(x,N)\leq\delta(M,N)\|x\|.$
Then we have
$\operatorname{dist}(Ax,BN)$ $\leq\|Ax-Bx\|+\operatorname{dist}(Bx,BN)$
$\leq\|Ax-Bx\|+\|B\|\operatorname{dist}(x,N)$
$\leq\|A-B\|\|x\|+\|B\|\|x\|\delta(M,N).$
Thus we obtain
$\delta(AM,BN)$ $=\sup_{x\in M\setminus\{0\}}\frac{\operatorname{dist}(Ax,BN)}{\|Ax\|}$
$\leq\|A^{-1}\|(\|A-B\|+\|B\|\delta(M,N)).$
We have the following key fact.
Lemma 2.21.
Let $(X,\omega)$ be a strong symplectic Banach space with a closed isotropic subspace $\lambda$. Then the following hold.
  • (a)
    The symplectic reduction $(\lambda^{\omega}/\lambda,\widetilde{\omega})$ is a strong symplectic Banach space.
  • (b)
    Assume that $\lambda$ is a maximal isotropic subspace with $\lambda\subsetneq\lambda^{\omega}$. Then we have $\gamma_{\lambda}>0$.
Proof.
(a) By [5, Lemma 1.2.6], we have $\lambda=\lambda^{\omega\omega}$. Let $f\in(\lambda^{\omega}/\lambda)^{*}$ be a bounded semilinear form. Then we have $f\pi_{\lambda}\in(\lambda^{\omega})^{*}$. By Hahn-Banach theorem, there is an $F\in X^{*}$ such that $F|_{\lambda^{\omega}}=f\pi_{\lambda}$. Since $(X,\omega)$ is a strong symplectic Banach space, there is an $x\in X$ such that $\omega(x,y)=Fy$ for all $y\in X$. Then we have
$\omega(x,y)=Fy=(f\pi_{\lambda})y=0$
for all $y\in\lambda$. Consequently, we have $x\in\lambda^{\omega}$. Then we obtain $f=L_{\tilde{\omega}}(x+\lambda)$. Since $f$ is arbitrary, we obtain that $(\lambda^{\omega}/\lambda,\widetilde{\omega})$ is a strong symplectic Banach space.
(b) By Proposition 2.17.c, there is an $h=\pm 1$ such that $\mathrm{i}h_{\lambda}\tilde{\omega}$ is positive definite on $\lambda^{\omega}/\lambda$. Then (39) follows from (a). ∎
Theorem 2.22.
Let $(X,\omega_{0})$ be a strong symplectic Banach space with a maximal isotropic subspace $\lambda$. Let $(X,\omega)$ be a symplectic Banach space with a closed isotropic subspace of $\mu$. We set
$a\ :$ $=\|L_{\omega_{0}}^{-1}\|\|L_{\omega_{0}}\|,$
$b(\omega)\ :$ $=\|L_{\omega_{0}}^{-1}\|\|L_{\omega}-L_{\omega_{0}}\|,$
$f(\mu,\omega)\ :$ $=b(\omega)+(a+b(\omega))\delta(\lambda,\mu),$
$g(\mu,\omega)\ :$ $=\left(|h_{\lambda}|\gamma_{\lambda}^{-1}\left(\|L_{\omega}-L_{\omega_{0}}\|+\|L_{\omega_{0}}\|f(\mu,\omega)(2+f(\mu,\omega))\right)\right)^{\frac{1}{2}}.$
Assume that there hold that
(42) $f(\mu,\omega)+g(\mu,\omega)<\frac{1-\delta(\lambda,\mu)}{1+\delta(\lambda,\mu)}.$
Then $\mu$ is a maximal isotropic subspace of the strong symplectic Banach space $(X,\omega)$ with $h_{\mu}\in\{h_{\lambda},0\}$.
Proof.
Since $L_{\omega_{0}}$ is a invertible skew-adjoint operator, by Proposition 2.11, $X$ is reflexive. Since $b(\omega)<1$, $L_{\omega}$ has bounded inverse. Then $(X,\omega)$ is a strong symplectic Banach space, and we have $\lambda^{\omega_{0}}=(L_{\omega_{0}}\lambda)^{\bot}$ and $\mu^{\omega}=(L_{\omega}\mu)^{\bot}$. By Lemma 2.21, $\gamma_{\lambda}>0$ holds if $h_{\lambda}\neq 0$.
By Lemma 2.20 and [13, Theorem IV.2.9], we have
$\delta(\mu^{\omega},\lambda^{\omega_{0}})\leq f(\mu,\omega).$
Then we have
$l(\mu,\omega)\leq g(\mu,\omega).$
By Theorem 1.1 and (42), $\mu$ is a Lagrangian (maximal isotropic) subspace of the strong symplectic Banach space $(X,\omega)$ with $h_{\mu}\in\{h_{\lambda},0\}$. ∎
Theorem 2.23.
Let $X$ be a Banach space with a continuous family of strong symplectic structure $\{\omega(s)\}_{s\in J}$, where $J$ is a connected topological space. Let $\{\lambda(s)\}_{s\in J}$ be a continuous family of closed subspace of $X$ such that $\lambda(s)\subset\lambda(s)^{\omega(s)}$ holds for each $s\in J$. Assume that there is an $s_{0}\in J$ such that $\lambda(s_{0})$ is a maximal isotropic subspace of $(X,\omega(s_{0}))$. Then $\lambda(s)$ is a maximal isotropic subspace of $(X,\omega(s))$ with $h_{\lambda(s)}=h_{\lambda(s_{0})}$ for each $s\in J$.
Proof.
By [5, Lemma 1.2.6], we have $\lambda(s)=\lambda(s)^{\omega(s)\omega(s)}$ for each $s\in J$. By Lemma 2.21, we have $\gamma_{\lambda(s_{0})}>0$ if $h_{\lambda(s_{0})}\neq 0$, and each $s\in J$ such that $\lambda(s)$ is a maximal isotropic subspace of $(X,\omega(s))$ satisfying $\gamma_{\lambda(s)}>0$ if $h_{\lambda(s)}\neq 0$. Then our result follows from Theorem 1.2. ∎
3. Stabilities with relative bounded perturbations
Although Theorem 1.1 is rather general, it is not very convenient for application since the definition of the gap $\delta(\lambda,\mu)$ is complicated. A less general but more convenient criterion is furnished by relatively bounded perturbations.
In this section we will prove similar stability results in [13, Sections V.4.1, V.4.2] in our setting. Firstly we introduce the notion of relative boundedness (cf. [13, Section IV.1.1], [8, Theorem III.7.4]).
Definition 3.1.
Let $X$, $Y$ be two Banach spaces and $T\colon X\leadsto Y$ be a linear relation. A linear relation $A\colon X\leadsto Y$ is called to be relative bounded with respect to $T$ or simply $T$-bounded if $\operatorname{dom}(T)\subset\operatorname{dom}(A)$, $\overline{T0}\supset A0$, and for some $p\geq 1$, there are non-negative constants $a_{p}$, $b_{p}$ such that
(43) $\inf\limits_{v\in Ax}\|v\|\leq\left(a_{p}^{p}\|x\|^{p}+b_{p}^{p}\|y\|^{p}\right)^{\frac{1}{p}}.$
for each $x\in\operatorname{dom}(T)$ and $y\in Tx$. The greatest lower bound $b_{0}$ of all possible constants $b_{p}$ in (43) will be called the relative bound of $A$ with respect to $T$ or simply the T-bound of $A$.
As in [13, Section V.4.1], the number $b_{0}$ does not depend on the choice of $p\geq 1$.
We have the following (cf. the proof of [13, Theorem V.4.4]). Recall [8, Definition II.5.4]:
Definition 3.2.
Let $X$, $Y$ be two Banach spaces. Let $T$ be a linear relation from $X$ to $Y$. $T$ is called closable if $\bar{T}$ is an extension of $T$.
Lemma 3.3.
Let $X$, $Y$ be two Banach spaces and $T\colon X\leadsto Y$ be a closable linear relation. Let $A\colon X\leadsto Y$ be a $T$-bounded linear relation such that (43) holds for some $p\geq 1$. Then $\bar{A}+T0\colon X\leadsto Y$ is a $\bar{T}$-bounded linear relation with the same constants $a_{p}$ and $b_{p}$, that is,
(44) $\inf\limits_{v\in\bar{A}x+T0}\|v\|\leq\left(a_{p}^{p}\|x\|^{p}+b_{p}^{p}\|y\|^{p}\right)^{\frac{1}{p}}.$
for each $x\in\operatorname{dom}(\bar{T})$ and $y\in\bar{T}x$. Moreover, we have
(45) $\overline{T+A}\supset\bar{T}+\bar{A}.$
Proof.
Since $T$ is closable, the set $T0=\bar{T}0$ is closed in $X$. Denote by $\pi\colon X\times Y\to(X\times Y)/(\{0\}\times T0)$ the natural projection along $\{0\}\times T0$. By the proof of [13, Theorem V.4.4], we have
$\|\pi(\bar{A})x\|$ $\leq\left(a_{p}^{p}\|x\|^{p}+b_{p}^{p}\|\pi(\bar{T})x\|^{p}\right)^{\frac{1}{p}},$
$\overline{\pi(T)+\pi(A)}$ $\supset\overline{\pi(T)}+\overline{\pi(A)}.$
Since $\overline{T0}\supset A0$, our lemma follows. ∎
We need a stability result for the closeness.
Proposition 3.4.
Let $X$, $Y$ be two Banach spaces. Let $T$, $S$ be linear relations from $X$ to $Y$ such that $S0=T0=W$, and
(46) $\inf\limits_{v\in(T-S)x}\|v\|\leq a\|x\|+b^{\prime}\inf\limits_{y\in Tx}\|y\|+b^{\prime\prime}\inf\limits_{y\in Sx}\|y\|$
for all $x\in\operatorname{dom}(T)=\operatorname{dom}(S)=D$, where $a$, $b^{\prime}$, $b$ are nonnegative constants and $b^{\prime}<1$, $b^{\prime\prime}<1$. Set $b\ :=\max\{b^{\prime},b^{\prime\prime}\}$ and $T(\kappa)\ :=T+\kappa(T-S)$ for all $\kappa\in[0,1]$. Then $S$ is closable if and only if $T$ is closable; in this case the closures of $T$ and $S$ have the same domain, and we have
(47) $\hat{\delta}(\overline{T(\kappa^{\prime})},\overline{T(\kappa)})\leq\frac{|\kappa^{\prime}-\kappa|(a^{2}+(b^{\prime}+b^{\prime\prime})^{2})^{\frac{1}{2}}}{1-b-|\kappa^{\prime}-\kappa|(b^{\prime}+b^{\prime\prime})}$
for all $\kappa^{\prime},\kappa\in[0,1]$ with $|\kappa^{\prime}-\kappa|(b^{\prime}+b^{\prime\prime})<1-b$. In particular $S$ is closed if and only if $T$ is.
Proof.
1. Denote by $\pi\colon X\times Y\to(X\times Y)/(\{0\}\times W)$ the natural projection along $\{0\}\times W$. Either $S$ or $T$ is closable will get that $W$ is closed. Then $X$ and $Y/W$ are Banach spaces, and $\pi(T)$ and $\pi(S)$ are linear operators from $X$ to $Y/W$ with the same domain $D$. By (46), we have
(48) $\|\pi(T-S)x\|\leq a\|x\|+b^{\prime}\|\pi(T)x\|+b^{\prime\prime}\|\pi(S)x\|$
for all $x\in D$. By [13, Theorem IV.1.3], $S$ is closable if and only if $T$ is closable; in this case the closures of $T$ and $S$ have the same domain. In particular $S$ is closed if and only if $T$ is.
2. Assume that both $T$ and $S$ are closable. By (48), we have
(49) $\|\pi(\bar{T}-\bar{S})x\|\leq a\|x\|+b^{\prime}\|\pi(\bar{T})x\|+b^{\prime\prime}\|\pi(\bar{S})x\|$
for all $x\in\operatorname{dom}(\bar{T})=\operatorname{dom}(\bar{S})$. Set $A\ :=\pi(\bar{T}-\bar{S})$. By [13, (IV.1.5)], we have
(50) $\|Ax\|\leq\frac{1}{1-b}\left(a\|x\|+(b^{\prime}+b^{\prime\prime})\|\pi(\overline{T(\kappa)})x\|\right)$
for all $x\in\operatorname{dom}(\bar{T})=\operatorname{dom}(\bar{S})$ and $\kappa\in[0,1]$. By [5, Lemma A.3.1.d] and [13, Theorem IV.2.14], we have
$\hat{\delta}(\overline{T(\kappa^{\prime})},\overline{T(\kappa)})$ $=\hat{\delta}(\pi(\overline{T(\kappa^{\prime})}),\pi(\overline{T(\kappa)}))$
$\leq\frac{|\kappa^{\prime}-\kappa|(a^{2}+(b^{\prime}+b^{\prime\prime})^{2})^{\frac{1}{2}}}{1-b-|\kappa^{\prime}-\kappa|(b^{\prime}+b^{\prime\prime})}$
for all $\kappa^{\prime},\kappa\in[0,1]$ with $|\kappa^{\prime}-\kappa|(b^{\prime}+b^{\prime\prime})<1-b$. ∎
We introduce the notions of essentially selfadjoint and essentially maximal symmetric as follows.
Definition 3.5.
Let $X$ and $Y$ be two Banach spaces over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{K}}$ be a nondegenerate sesquilinear form. Let $\omega$ be the associated symplectic form of $\Omega$ defined by (14). Let $A\colon X\leadsto Y$ be a linear relation. We call $A$ essentially selfadjoint (essentially maximal symmetric, respectively) with respect to $\Omega$ if $A$ is closable and $\bar{A}$ is selfadjoint (maximal isotropic,respectively).
Then we have the following similar result of [13, Theorem V.4.5]. Note that the special case of Hilbert space situation [13, Theorems V.4.3 and V.4.4] is due to F. Rellich [14].
Theorem 3.6.
Let $X$ and $Y$ be two Banach spaces over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{K}}$ be a nondegenerate sesquilinear form with $L_{\Omega}^{-1}\in{\mathcal{B}}(Y^{*},X)$. Let $T$, $S$ be two symmetric relations with respect to $\Omega$ such that $\operatorname{dom}(T)=\operatorname{dom}(S)=D$, $T0=S0$ and
(51) $\inf\limits_{v\in(T-S)x}\|v\|\leq a\|x\|+b(\inf\limits_{y\in Tx}\|y\|+\inf\limits_{y\in Sx}\|y\|)$
holds for all $x\in D$, where $a$, $b$ are nonnegative constants and $b<1$. Then $S$ is essentially selfadjoint (or maximal symmetric) with respect to $\Omega$ if and only if $T$ is; in this case, $\bar{S}$ and $\bar{T}$ have the same domain. In particular, $S$ is selfadjoint (or maximal symmetric) with respect to $\Omega$ if and only if $T$ is.
Proof.
Since $L_{\Omega}^{-1}\in{\mathcal{B}}(Y^{*},X)$, the associated symplectic form of $\Omega$ defined by (14) is strong. By Proposition 3.4, we can replace $S$ and $T$ by $\bar{S}$ and $\bar{T}$. Then our results follows from Proposition 3.4 and Theorem 2.23. ∎
4. Skew-adjoint relations between real Banach spaces
4.1. Skew-adjoint relations between real Hilbert spaces
We denote by $V_{{\mathbb{C}}}\ :=V\otimes_{{\mathbb{R}}}{\mathbb{C}}$ for each real vector space $V$. Let $(X,\langle\cdot,\cdot\rangle)$ be a real Hilbert space. Then we have $\Omega_{X}(x,y)=\langle x,y\rangle$ for each $x,y\in X$. Denote by $I_{X}$ the identity map on $X$, and
$J_{X^{2}}\>=\left(\begin{array}[]{cc}0&-I_{X}\\ I_{X}&0\end{array}\right).$
Then we have $\omega(u,v)=\langle J_{X^{2}}u,v\rangle$ for all $u,v\in X^{2}$, where $\omega$ is the symplectic structure of $X^{2}$ associated with $\Omega_{X}$. We denote by $\mathrm{U}(X_{\mathbb{C}})$ the unitary group on $X_{\mathbb{C}}$ and $\mathrm{O}(X)$ the orthogonal group on $X$. We have the following structure of ${\mathcal{R}}^{sk}(X)$.
Proposition 4.1.
Let $(X,Q=\langle\cdot,\cdot\rangle)$ be a real Hilbert space and $Y$ be a real linear space. Let $\Omega\colon X\times Y$ be a bilinear form such that $R_{\Omega}\colon Y\to X^{*}$ is a real linear isomorphism. Then we have a homemorphism $\varphi_{\Omega}\colon\mathrm{U}(X_{\mathbb{C}})\to{\mathcal{R}}^{sk}(\Omega_{\mathbb{C}})$ defined by
(52) $\varphi_{\Omega}(U)=R_{\Omega}^{-1}R_{Q}\circ(I_{X_{C}}-U)\circ(I_{X_{C}}+U)^{-1}$
for all $U\in\mathrm{U}(X_{C})$ such that $\varphi_{\Omega}(\mathrm{O}(X))=\{A_{\mathbb{C}};\;A\in{\mathcal{R}}^{sk}(\Omega)\}$. For each $T=\varphi_{\Omega}(U)\in{\mathcal{R}}^{sk}(\Omega_{\mathbb{C}})$, we have $\ker_{\mathbb{C}}T=\ker_{\mathbb{C}}(U-I)$ and $T_{\mathbb{C}}0=R_{\Omega}^{-1}R_{Q}\ker_{\mathbb{C}}(U+I)$.
Proof.
1. Denote by $X^{\pm}\ :=\{(x,\pm\mathrm{i}x);\;x\in X_{\mathbb{C}}\}$, Then we have $J_{X^{2}}u=\mp\mathrm{i}u$ for all $u\in X^{\pm}$. By [4, Proposition 2] we have
${\mathcal{R}}^{sa}(X_{\mathbb{C}})$ =$=\{\{(x,\mathrm{i}x)+(Ux,-\mathrm{i}Ux);\;x\in X_{\mathbb{C}}\};\;U\in\mathrm{U}(X_{\mathbb{C}})\}$
=$=\{\mathrm{i}(I_{X_{C}}-U)\circ(I_{X_{C}}+U)^{-1};\;U\in\mathrm{U}(X_{\mathbb{C}})\}.$
Then we have
${\mathcal{R}}^{sk}(X_{\mathbb{C}})$ =$=\{\{(x,x)+(Ux,-Ux);\;x\in X_{\mathbb{C}}\};\;U\in\mathrm{U}(X_{\mathbb{C}})\}$
=$=\{(I_{X_{C}}-U)\circ(I_{X_{C}}+U)^{-1};\;U\in\mathrm{U}(X_{\mathbb{C}})\}.$
Then $\varphi_{\Omega_{X}}$ defined by (52) is a homemorphism and $\varphi_{\Omega_{X}}(\mathrm{O}(X))\subset\{A_{\mathbb{C}};\;A\in{\mathcal{R}}^{sk}(X)\}$.
2. Let $A\in\{A_{\mathbb{C}};\;A\in{\mathcal{R}}^{sk}(X)\}$. By Step 1, there is a $U=C+\mathrm{i}D\in\mathrm{U}(X_{\mathbb{C}})$, $C,D\in{\mathcal{B}}(X)$ such that
$A_{\mathbb{C}}$ =$=\{(x,x)+(Ux,-Ux);\;x\in X_{\mathbb{C}}\}$
=$=\{(u(x,y),v(x,y));\;x,y\in X\}.$
where $u(x,y)=x+Cx-Dy+\mathrm{i}(y+Cy+Dx)$ and $v(x,y)=x-Cx+Dy+\mathrm{i}(y-Cy-Dx)$. From this we obtain
$A$ =$=\{(x+Cx-Dy,x-Cx+Dy);\;x,y\in X\}$
=$=\{(y+Cy+Dx,y-Cy-Dx);\;x,y\in X\}$
$\subset A_{\mathbb{C}}.$
We conclude that $Cx-Dy=Ux$ for all $x,y\in X$. Then we have $D=0$ and $U=C\in\mathrm{O}(X)$.
3. Since
${\mathcal{R}}^{sk}(X_{\mathbb{C}})=\{\{(x,R_{Q}^{-1}R_{\Omega}y);\;(x,y)\in T\};\;T\in{\mathcal{R}}^{sk}(\Omega_{\mathbb{C}})\},$
by Step 1 and Step 2, the map $\varphi_{\Omega}$ defined by (52) satisfies the desired properties. ∎
4.2. Stability theorem of mod 2 index
In this subsection we prove Theorem 1.3. We need the following fact.
Lemma 4.2.
Let $X$, $Y$ be two finite dimensional vector space over ${\mathbb{K}}$. Let $\Omega\colon X\times Y\to{\mathbb{K}}$ be a nondegenerate sesquilinear form. Let $T$ be a selfadjoint (skew-adjoint, respectively) linear relation with respect to $\Omega$. Let $Q_{T}$ be the associated form of $T$ defined by Definition 2.4. Then we have
(53) $\ker T=\ker Q_{T}.$
Proof.
Clearly $Q$ is well-defined and $\ker T\subset\ker Q_{T}$. Since $T$ is selfadjoint (skew-adjoint, respectively), we have $\dim T=\dim X=\dim Y$. Since $R_{\Omega}$ is a linear isomorphism, we have
$\dim\ker Q_{T}=$ $\dim(\operatorname{im}T)^{\Omega,l}=\dim Y/\operatorname{im}T$
$=$ $\dim T-\dim\operatorname{im}T=\dim\ker T.$
Since $\ker T\subset\ker Q_{T}$, we have $\ker T=\ker Q_{T}$. ∎
Proof of Theorem 1.3.
We divide the proof into five steps. We use the same symbol to denote the complexified sesquilinear forms of the corresponding real bilinear form.
1. Since $T\colon X\leadsto Y$ is a skew-adjoint linear Fredholm relation with respect to $\Omega_{0}$ with index 00, there is a real linear subspace $V$ of $Y$ such that
(54) $Y=\operatorname{im}T\oplus V,\quad\dim\ker T=\dim V<+\infty.$
2. Assume that $\dim X<+\infty$. Define $Q_{T}\colon\operatorname{dom}(T)\times\operatorname{dom}(T)\to{\mathbb{R}}$ by $Q_{T}(x,y)\ :=\Omega_{0}(x,Ty)$ for all $x,y\in\operatorname{dom}(T)$. Since $T$ is skew-adjoint, we have $\dim T=\dim X=\dim Y$. By Witt decomposition theorem we have
(55) $\dim\operatorname{dom}(T)=\dim X-\dim T0=2m^{-}(\mathrm{i}Q_{T})+\dim\ker Q_{T}.$
By Lemma 4.2 we have $\dim\ker Q_{T}=\dim\ker T$.
3. Set $Z\ :=X_{{\mathbb{C}}}\times Y_{{\mathbb{C}}}$. We denote by $\omega_{0}$ and $\omega$ the associated symplectic structure of $\Omega_{0}$ and $\Omega$ on $Z$ respectively. We set
$\alpha\ :=X_{{\mathbb{C}}}\times\{0\},\quad\beta\ :=\{0\}\times V_{{\mathbb{C}}},\quad\gamma\ :=(\mathrm{i}S)_{{\mathbb{C}}},\quad\lambda\ :=(\mathrm{i}T)_{{\mathbb{C}}}.$
Then $\alpha$ is a Lagrangian subspace of $(Z,\omega_{0})$ and $(Z,\omega)$, $\beta$ is a finite dimensional isotropic subspace of $(Z,\omega_{0})$ and $(Z,\omega)$, $\gamma$ is an isotropic subspace of $(Z,\omega)$, and $\lambda$ is a Lagrangian subspace of $(Z,\omega_{0})$. By [5, Proposition 1.3.3] we have
$Z=\alpha\cap\beta^{\omega_{0}}\oplus\beta\oplus\lambda=\alpha+\beta^{\omega_{0}}.$
We have $(\alpha\cap\beta^{\omega})^{\omega}\supset\alpha+\beta$. By [5, Lemma 1.1.2], we have
$\dim_{{\mathbb{C}}}(\alpha\cap\beta^{\omega})^{\omega}/\alpha\leq$ $\dim_{\mathbb{C}}\alpha/(\alpha\cap\beta^{\omega})$
$=$ $\dim_{\mathbb{C}}X_{\mathbb{C}}/V_{\mathbb{C}}^{\Omega,l}$
$=$ $\dim_{\mathbb{C}}V_{\mathbb{C}}=\dim_{\mathbb{C}}\beta.$
Then we have
$(\alpha\cap\beta^{\omega})^{\omega}=\alpha+\beta.$
By [13, Lemma IV.2.9] and [5, Lemma 3.1.1], there is a $\delta>0$ such that for each closed skew-symmtric linear relation $S$ with respect to $\Omega$ with $\hat{\delta}(S,T)+\|L_{\Omega}-L_{\Omega_{0}}\|<\delta$, we have
$\{0\}$ $=\beta\cap\gamma,$
$Z$ $=\alpha\cap\beta^{\omega}+\beta+\gamma=\alpha+\beta^{\omega},$
$\{0\}$ $=Z^{\omega}=(\alpha+\beta+\gamma)^{\omega}=\alpha\cap\beta^{\omega}\cap\gamma,$
By [13, Theorem IV.4.30, Problem IV.4.6] we have
$0=\operatorname{index}T=\operatorname{index}S=\operatorname{index}_{\mathbb{C}}(\alpha+\beta,\gamma)-\dim_{\mathbb{C}}\beta.$
By Lemma 2.10, $S$ is skew-adjoint with respect to $\Omega$.
4. We define the sesquilinear form $Q$ on $\alpha\cap(\beta+\gamma)$ by $Q(x_{1},x_{2})=\omega(x_{1},y_{2})$ if $x_{j}\in\alpha$, $y_{j}\in\beta$, $x_{j}+y_{j}\in\gamma$, $j=1,2$. By [16, §3.1], the form $Q$ is a well-defined symmetric form.
Let $\alpha_{1}$ be a complex linear subspace $\alpha$ with $\dim_{\mathbb{C}}\alpha/\alpha_{1}<+\infty$. By [13, Problem IV.4.6] we have
$\dim_{\mathbb{C}}(\gamma\cap(\alpha_{1}+\beta))$ $=\operatorname{index}_{\mathbb{C}}(\alpha_{1}+\beta,\gamma)+\dim_{\mathbb{C}}Z/(\alpha_{1}+\beta+\gamma)$
$=\operatorname{index}_{\mathbb{C}}(\alpha_{1},\gamma)+\dim\beta+\dim_{\mathbb{C}}Z/(\alpha_{1}+\beta+\gamma)$
$=\dim_{\mathbb{C}}\beta+\dim_{\mathbb{C}}Z/(\alpha_{1}+\beta+\gamma)-\dim_{\mathbb{C}}\alpha/\alpha_{1}.$
Since $\alpha_{1}\cap\beta=\beta\cap\gamma=\{0\}$, there is a linear isomorphism $\phi\colon\gamma\cap(\alpha_{1}+\beta)\to\alpha_{1}\cap(\beta+\gamma)$ defined by $z-\phi(z)\in\beta$. Then we have
(56) $\begin{split}\dim_{\mathbb{C}}(\alpha_{1}\cap(\beta+\gamma))=\dim\beta+\dim_{\mathbb{C}}Z/(\alpha_{1}+\beta+\gamma)-\dim_{\mathbb{C}}\alpha/\alpha_{1}.\end{split}$
Since $\alpha+\beta+\gamma=Z$, we have
(57) $\begin{split}\dim_{\mathbb{C}}(\alpha\cap(\beta+\gamma))=\dim_{\mathbb{C}}\beta=\dim\ker T.\end{split}$
We have $(\alpha+\beta)^{\omega}=\alpha\cap\beta^{\omega}$ and $(\alpha+\beta)^{\omega\omega}=(\alpha\cap\beta^{\omega})^{\omega}=\alpha+\beta$. Since $Z=\alpha+\beta^{\omega}$, by [5, Lemma 1.3.2] we have
$\dim_{\mathbb{C}}\alpha/(\alpha\cap\beta^{\omega})=\dim_{\mathbb{C}}Z/\beta^{\omega}=\dim_{\mathbb{C}}\beta.$
By [13, Problem IV.4.6] we have
$\operatorname{index}(\alpha+$ $\beta,\gamma)+\operatorname{index}((\alpha+\beta)^{\omega},\gamma)$
$=$ $\operatorname{index}(\alpha,\gamma)+\dim_{\mathbb{C}}(\alpha+\beta))/\alpha+\operatorname{index}(\alpha,\gamma)-\dim_{\mathbb{C}}\alpha/(\alpha+\beta)^{\omega}$
$=$ $0\displaystyle 2\operatorname{index}(\alpha,\gamma)=$ 0 .
By [5, Corollary 1.2.12] we have
$(\gamma\cap(\alpha+\beta))^{\omega}=\gamma+(\alpha+\beta)^{\omega}=\gamma+\alpha\cap\beta^{\omega}.$
By [16, Lemma 3.3] and [5, Lamma A.1.1] we have
$\ker Q$ $=\alpha\cap(\beta\cap(\gamma\cap(\alpha+\beta))^{\omega}+\gamma)$
$=\alpha\cap(\beta\cap(\gamma+\alpha\cap\beta^{\omega})+\gamma)$
$=\alpha\cap(\beta+\gamma)\cap(\gamma+\alpha\cap\beta^{\omega})$
$=(\beta+\gamma)\cap(\alpha\cap\gamma+\alpha\cap\beta^{\omega}).$
Since $\alpha\cap\beta^{\omega}+\beta+\gamma=Z$ and $\alpha\cap\beta^{\omega}\cap\gamma=\{0\}$, by (56) and [13, Problem IV.4.6] we have
$\dim_{\mathbb{C}}\ker Q=$ $\dim_{\mathbb{C}}\beta-\dim_{\mathbb{C}}\alpha/(\alpha\cap\gamma+\alpha\cap\beta^{\omega})$
$=$ $\dim_{\mathbb{C}}\beta-\dim_{\mathbb{C}}\alpha/(\alpha\cap\beta^{\omega})+\dim_{\mathbb{C}}(\alpha\cap\gamma+\alpha\cap\beta^{\omega})/(\alpha\cap\beta^{\omega})$
$=$ $\dim_{\mathbb{C}}(\alpha\cap\gamma)/(\alpha\cap\beta^{\omega}\cap\gamma)$
$=$ $\dim_{\mathbb{C}}(\alpha\cap\gamma)=\dim\ker S.$
5. We have $\alpha\cap(\beta+\gamma)=S_{\mathbb{C}}^{-1}V_{{\mathbb{C}}}\times\{0\}$. For all $x_{1},x_{2}\in S_{\mathbb{C}}^{-1}V_{{\mathbb{C}}}$, we take a $y_{1}\in V_{{\mathbb{C}}}\cap S_{\mathbb{C}}x_{1}$ and a $y_{2}\in V_{{\mathbb{C}}}\cap S_{\mathbb{C}}x_{2}$. Then we have
$Q((x_{1},0),(x_{2},0))$ $=\omega((x_{1},0),(0,\mathrm{i}y_{2}))$
$=\Omega(x_{1},\mathrm{i}y_{2}).$
Note that for $x_{1},x_{2}\in S^{-1}V$, we can take a $y_{1}\in V\cap Sx_{1}$ and a $y_{2}\in V\cap Sx_{2}$. Then $\mathrm{i}Q$ is a real skew symmetric form on the real vector space $S^{-1}V$. By Step 1, Step 4, (55) and (57) we have
$\dim\ker T=$ $\dim_{\mathbb{C}}(\alpha\cap(\beta+\gamma))=\dim\ker\mathrm{i}Q|_{S^{-1}V}$
$=$ $\dim_{\mathbb{C}}\ker_{{\mathbb{C}}}Q=\dim\ker S\mod 2.$
4.3. Path components
In this subsection, we study the path components of the set of skew-adjoint Fredholm relations between real Banach spaces with index 00 and prove Theorem 1.4.
Let $X$ and $Y$ be two Banach space, and $\Omega\colon X\times Y\to{\mathbb{K}}$ be a bounded nondegenerate sesquilinear form. For $k\in{\mathbb{N}}$, we define
(58) ${\mathcal{R}}_{k,0}^{sk}(\Omega)\ :$ =$=\{T\in{\mathcal{R}}_{0}^{sk}(\Omega);\;\dim\ker T=k\}.$
We need the following lemmas.
Lemma 4.3.
Let $X$ and $Y$ be two vector space over ${\mathbb{K}}$, and $\Omega\colon X\times Y$ be a nondegenerate sesquilinear form. Then ${\mathcal{R}}_{0,0}^{sk}(\Omega)$ is an affine space over ${\mathbb{R}}$.
Proof.
Since ${\mathcal{R}}_{0,0}^{sk}(\Omega)=\{T\in{\mathcal{R}}^{sk}(\Omega);\;T^{-1}\in\operatorname{Hom}(Y,X)\}$, the set ${\mathcal{R}}_{0,0}^{sk}(\Omega)$ is an affine space over ${\mathbb{R}}$. ∎
Lemma 4.4.
Let $X$ and $Y$ be two vector space over ${\mathbb{K}}$, and $\Omega\colon X\times Y$ be a nondegenerate sesquilinear form. Then the following hold.
  • (a)
    Let $X_{0}$, $X_{1}$ be linear subspaces of $X$ and $Y_{0}$, $Y_{1}$ be linear subspaces of $Y$ such that
    (59) $\begin{split}&X=X_{0}+X_{1},\quad Y=Y_{0}+Y_{1},\\ &\Omega(x,y)=0\quad\forall(x,y)\in(X_{0}\times Y_{1})\cup(X_{1}\times Y_{0}).\end{split}$
    Then $\Omega|_{X_{0}\times Y_{0}}$ and $\Omega|_{X_{1}\times Y_{1}}$ are nondegenerate, and we have
    (60) $X=X_{0}\oplus X_{1},\quad Y=Y_{0}\oplus Y_{1},$
    (61) $X_{0}=Y_{1}^{\Omega,l},\quad X_{1}=Y_{0}^{\Omega,l},\quad Y_{0}=X_{1}^{\Omega,r},\quad Y_{1}=X_{0}^{\Omega,r},$
    (62) ${\mathcal{R}}^{sk}(\Omega)\supset\{T_{0}\oplus T_{1};\;T_{0}\in{\mathcal{R}}^{sk}(\Omega|_{X_{0}\times Y_{0}}),T_{1}\in{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}})\}.$
  • (b)
    Let $T\colon X\leadsto Y$ be a skew-adjoint linear Fredholm relation with index 00. We denote by $X_{0}\ :=\ker T$ and $Y_{1}\ :=\operatorname{im}T$. Let $Y_{0}$ be a linear subspace of $Y$ such that $Y=Y_{0}\oplus Y_{1}$. We denote by $X_{1}\ :=Y_{0}^{\Omega,l}$. Denote by $Q_{T}$ the associated form of $T$ defined by Definition 2.4.c. Then (59) holds, and we have
    (63) $T$ $=T_{0}\oplus T_{1},\quad Q_{T}=Q_{T_{0}}\oplus Q_{T_{1}},\quad\ker T=\ker Q_{T},$
    where
    (64) $T_{0}\ :$ $=X_{0}\times\{0\}\in{\mathcal{R}}^{sk}(\Omega|_{X_{0}\times Y_{0}}),$
    (65) $T_{1}\ :$ $=T\cap(X_{1}\times Y_{1})\in{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}}).$
Proof.
(a) Since $Y=Y_{0}+Y_{1}$ and $X_{0}\subset Y_{1}^{\Omega,l}$ hold, we have
$X_{0}\cap Y_{0}^{\Omega,l}=X_{0}\cap Y_{1}^{\Omega,l}\cap Y_{0}^{\Omega,l}=X_{0}\cap(Y_{0}+Y_{1})^{\Omega,l}=\{0\}.$
Similarly we have $Y_{0}\cap X_{0}^{\Omega,r}=\{0\}$. Then $\Omega|_{X_{0}\times Y_{0}}$ is nondegenerate. Similarly $\Omega|_{X_{1}\times Y_{1}}$ is nondegenerate.
Since $Y=Y_{0}+Y_{1}$, $X_{0}\subset Y_{1}^{\Omega,l}$ and and $X_{1}\subset Y_{0}^{\Omega,l}$ hold, we have
$X_{0}\cap X_{1}\subset Y_{0}^{\Omega,l}\cap Y_{1}^{\Omega,l}=(Y_{0}+Y_{1})^{\Omega,l}=\{0\}.$
Then we have $X=X_{0}\oplus X_{1}$. Similarly, we have $Y=Y_{0}\oplus Y_{1}$.
Since $Y_{1}^{\Omega,l}\supset X_{0}$ and $Y_{1}^{\Omega,l}\cap X_{1}=\{0\}$, by [5, Lemma A.1.1] we have
$Y_{1}^{\Omega,l}=Y_{1}^{\Omega,l}\cap(X_{0}+X_{1})=X_{0}+Y_{1}^{\Omega,l}\cap X_{1}=X_{0}.$
Similarly we get the other three identities of (61).
The equation (62) is clear.
(b) By [5, Lemma 1.1.2.c] we have $\dim X_{0}=\dim Y_{0}=\dim X/X_{1}$. Since $\mathrm{i}T$ is maximal symmetric, by Proposition 2.9 we have $X_{0}=Y_{1}^{\Omega,l}$. Since $Y=Y_{0}\oplus Y_{1}$, we have $X_{0}\cap X_{1}=(Y_{0}+Y_{1})^{\Omega,l}=\{0\}$. Then we have $X=X_{0}\oplus X_{1}$. Then (59) holds. By (a), we have $T=T_{0}\oplus T_{1}$, where $T_{0}$ and $T_{1}$ are defined by (64) and (64) respectively. Then we have $Q_{T}=Q_{T_{0}}\oplus Q_{T_{1}}$. Since $\operatorname{im}T_{1}=Y_{1}$, we have $\ker Q_{T_{1}}=\{0\}$. Then we have
$\ker Q_{T}=\ker Q_{T_{0}}\oplus\ker Q_{T_{1}}=X_{0}=\ker T.$
Remark 4.5.
Lemma 4.4 holds for the selfadjoint case.
Lemma 4.6.
Let $X$ and $Y$ be two Banach space over ${\mathbb{K}}$, and $\Omega\colon X\times Y\to{\mathbb{K}}$ be a bounded nondegenerate sesquilinear form. Denote by $G(k,X)$ the Grassmanian of the linear subspace of $X$ of dimension $k$. We define the map $\pi\colon{\mathcal{R}}_{k,0}^{sk}(\Omega)\to G(k,X)$ by
(66) $\pi(T)=\ker T.$
Then $({\mathcal{R}}_{k,0}^{sk}(\Omega),\pi)$ is a fiber bundle with fiber $\pi^{-1}(X_{0})\simeq{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}})$ at $X_{0}$, where the notations $X_{i}$ and $Y_{i}$ with $i=0,1$ is defined in Lemma 4.4.a.
Proof.
Set $G(X,X_{1})\ :=\{M\in{\mathcal{S}}(X);X=M\oplus X_{1}\}$ By [5, (A.10)], the set $G(X,X_{1})$ is an open affine space and we have
(67) $G(X,X_{1})=\{\operatorname{graph}(A);A\in{\mathcal{B}}(X_{0},X_{1})\}.$
. By Lemma 4.4, we have $\pi^{-1}(X_{0})\simeq{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}})$.
Let $M=\operatorname{graph}(A)\in G(X,X_{1})$ be a linear subspace transversal to $X_{1}$ with $A\in{\mathcal{B}}(X_{0},X_{1})$. Then we have
$Y_{0}\cap M^{\Omega,r}=X_{1}^{\Omega,r}\cap M^{\Omega,r}=(X_{1}+M)^{\Omega,r}=\{0\}.$
By [5, Lemma 1.1.2.c], we have
$\dim Y/M^{\Omega,r}=\dim M=\dim X_{0}=\dim Y_{0}.$
Then we have $Y=Y_{0}\oplus M^{\Omega,r}$.
By [5, (A.10)], we have $M^{\Omega,r}=\operatorname{graph}(B)$ for a uniquely determined $B\in{\mathcal{B}}(Y_{1},Y_{0})$. Then we have
(68) $\Omega(x_{0},By_{1})+\Omega(Ax_{0},y_{1})=\Omega(x_{0}+Ax_{0},By_{1}+y_{1})=0$
for each $(x_{0},y_{1})\in X_{0}\times Y_{1}$.
We have
$S_{1}\in$ ${\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times M^{\Omega,r}})$
$\Leftrightarrow$ $S_{1}=(B+I_{Y_{1}})\circ T_{1},T_{1}\subset X_{1}\times Y_{1},S_{1}\oplus(X_{1}\times\{0\})=X_{1}\times M^{\Omega,r},$
$\Omega(x_{1},By_{2}+y_{2})=-\overline{\Omega(x_{2},By_{1}+y_{1})}\text{ for all }(x_{i},y_{i})\in T_{1},i=1,2$
$\Leftrightarrow$ $S_{1}=(B+I_{Y_{1}})\circ T_{1},T_{1}\subset X_{1}\times Y_{1},T_{1}\oplus(X_{1}\times\{0\})=X_{1}\times Y_{1},$
$\Omega(x_{1},y_{2})=-\overline{\Omega(x_{2},y_{1})}\text{ for all}(x_{i},y_{i})\in T_{1},i=1,2$
$\Leftrightarrow$ $S_{1}=(B+I_{Y_{1}})\circ T_{1},T_{1}\in{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}}).$
With the notations above, we define $\varphi_{X_{1},Y_{1}}:\pi^{-1}(G(X,X_{1}))\to G(X,X_{1})\times{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}})$ by
(69) $\varphi_{X_{1},Y_{1}}((M\times\{0\})\oplus((B+I_{Y_{1}})\circ T_{1}))=(M,T_{1}).$
Then $\varphi_{X_{1},Y_{1}}$ is a homemorphism. By definition, $({\mathcal{R}}_{k,0}^{sk}(\Omega),\pi)$ is a fiber bundle with fiber $\pi^{-1}(X_{0})\simeq{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}})$ at $X_{0}$. ∎
Corollary 4.7.
Let $X$ and $Y$ be two Banach space over ${\mathbb{K}}$, and $\Omega\colon X\times Y\to{\mathbb{K}}$ be a bounded nondegenerate sesquilinear form. Then the space ${\mathcal{R}}_{k,0}^{sk}(\Omega)$ is path connected for each nonnegative integer $k$.
Proof.
By Lemma [5, A.4.4], the space $G(k,X)$ is path connected. By Lemma 4.3, the space ${\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}})$ is path connected. By Lemma 4.6, the space ${\mathcal{R}}_{k,0}^{sk}(\Omega)$ is path connected. ∎
Proof of Theorem 1.4.
Let $a\in\{0,1\}$. By Lemma 4.4.a and Proposition 4.1, the set ${\mathcal{R}}_{a;0}^{sk}(\Omega)$ is nonempty. Take a $T\in{\mathcal{R}}_{a;0}^{sk}(\Omega)$ and use the notations in Lemma 4.4. By Lemma 4.4 and Proposition 4.1, there is a path $c_{1}\colon[0,1]\to\{S_{0}\oplus S_{1};\;S_{0}\in{\mathcal{R}}^{sk}(\Omega|_{X_{0}\times Y_{0}}),S_{1}\in{\mathcal{R}}_{0,0}^{sk}(\Omega|_{X_{1}\times Y_{1}})\}$ with $c_{0}(0)=T$ and $c_{0}(1)\in{\mathcal{R}}_{a,0}^{sk}(\Omega)$. By Corollary 4.7, the space ${\mathcal{R}}_{a,0}^{sk}(\Omega)$ is path connected. Then the space ${\mathcal{R}}_{a;0}^{sk}(\Omega)$ is path connected. By Theorem 1.3, there is no path joining ${\mathcal{R}}_{0;0}^{sk}(\Omega)$ and ${\mathcal{R}}_{1;0}^{sk}(\Omega)$. Then the space ${\mathcal{R}}^{sk}(\Omega)$ has eactly two path components ${\mathcal{R}}_{0;0}^{sk}(\Omega)$ and ${\mathcal{R}}_{1;0}^{sk}(\Omega)$. ∎